### v1

# SupplementaryMethods Ultra-high dose rate 6 MeV electron irradiation generates stable [1-13C]alanine radicals suitable for medical imaging with dissolution Dynamic Nuclear Polarisation

Abstract
Dissolution Dynamic Nuclear Polarisation (dDNP) is an experimental technique that increases the sensitivity of magnetic resonance experiments by more than a factor of , permitting isotopically-labelled molecules to be transiently visible in MRI scans with their biochemical fates spatially resolvable over time following injection into a patient. dDNP requires a source of unpaired electrons to be in contact with the isotope-labelled nuclei, cooled to temperatures close to absolute zero, and spin-pumped into a given state by microwave irradiation. At present, these electrons are typically provided by chemical radicals which require removal by filtration prior to injection into humans. Alternative sources include UV irradiation, requiring storing samples in liquid nitrogen, or cobalt-60 gamma irradiation, which requires days and generates polarisation two to three orders of magnitude lower than chemical radicals. In this study, we present ultra-high dose rate electron beam irradiation as a novel alternative for generating non-persistent radicals in glycerol/alanine mixtures. These radicals are stable for months at room temperature, are present at concentrations dependent on irradiation dose, and generate comparable nuclear polarisation to the typically used trityl radicals (20%) through a novel mechanism. The process of their generation inherently sterilises samples, and they enable the imaging of alanine metabolism in vivo using dDNP. This new method of generating radicals for dDNP offers the potential to report on relevant biological processes while being translatable to the clinic.
1 Introduction
Hyperpolarised Magnetic Resonance Imaging (HP MRI) is a molecular imaging technique that is predominantly used to monitor metabolic fluxes in real-time by transiently making a labelled molecular probe visible to magnetic resonance experiments such as MRI. Over the past two decades, its pre-clinical use has spanned multiple fields of basic scientific and medical research, with 2013 marking the start of its validation for various clinical applications, including prostate cancer[1], traumatic brain injury[2], and ischaemic heart disease,[3] amongst many others.[4] To perform HP MRI, a hyperpolariser is required – a device that transiently enhances the MR signal of stable labelled isotopes by several orders of magnitude. One method to achieve this is through dissolution Dynamic Nuclear Polarisation (dDNP), which employs cryogenic temperatures, high magnetic field strengths ( T), and microwave irradiation. A typical sample for dDNP comprises a metabolic substrate mixed with a chemical radical species and optionally also a glassing agent such as glycerol. At low ( K) temperature, the unpaired electron spins present in the radical become thermally polarised. When irradiated with microwaves, a population inversion of these electronic spins can be created, and they are driven to relax, primarily through dipolar interactions, and ultimately transfer their polarisation to nearby nuclei. The net process behind DNP therefore is that nuclear polarisation above that at thermal equilibrium is created. Once the desired nuclear signal enhancement is achieved, a heated and pressurised solvent is introduced to rapidly melt the solid-state sample and convert it into a form suitable for injection into a living system. Following injection, the enzymatic conversion of the hyperpolarised substrate into other metabolic intermediates can be temporally and spatially monitored in vivo through magnetic resonance experiments such as MRI.
Hyperpolarised pyruvate has become the most widely used substrate in HP MRI due to its molecular utility and importance in biochemistry, and the high degree of polarisation () that can be obtained with dDNP.[5] However, in clinical settings, its preparation is complicated by the relative toxicity of the chemical radicals used. This requires either (1) chemical engineering to produce a soluble trityl radical that can be filtered out – the approach taken in all clinical trials to date, as the LD50 of trityl radicals is approximately [6] and they are present in samples at concentrations above this – or (2) photon irradiation to generate transient radicals that serve as the electron source necessary for dDNP and quench rapidly upon dissolution.[7] The key advantages of the latter approach are the elimination of the filtration step, enabling faster and more efficient use of the sample, and the extended duration of the enhanced MR signal (i.e. a lengthened relaxation time), as it avoids the presence of unpaired electrons that act as relaxation centres following DNP[8]. To date, both ultraviolet[9] and gamma irradiation[10] have shown potential.
However, regardless of whether the electron source is chemical or photon-based, there are substantial technical limitations on the ability of clinical sites to ship and receive compounds for hyperpolarisation with the radicals in situ. For instance, the chemical instability of the trityl radical — the most commonly used chemical radical — renders centralised production of pre-filled fluid paths for global distribution infeasible; fluid paths are transported typically at K. In contrast, while low-energy gamma-generated radicals are reported to be stable at room temperature for several months to produce a single sample, irradiation can take days to months, and reported nuclear enhancement factors are on the order of and not the obtained with chemical radicals.[11] DNP with UV-generated radicals is only possible if samples are optically transparent (e.g. pyruvic acid) and require low-temperature (cryogenic) transportation[7], which incurs a significant expense and is not routinely performed as precise control of and is required over many orders of magnitude.[12] Ultimately, sites conducting human dDNP experiments generally require local access to a pharmacy equipped with sterile compounding facilities to comply with the high degree of sterility assurance required. This is complex from a regulatory view,[4] and is considered a significant financial burden that may ultimately limit the accessibility of the technique.
Here, we present an alternative approach for generating sterilised samples for clinical HP MRI that are suitable for convenient transport and centralised manufacture, potentially offering an alternative for sites that do not have access to a pharmacy. Using an ultra-high dose rate 6 MeV electron linear accelerator (linac), we demonstrate the creation of endogenous radicals within polycrystalline alanine, the concentration of which can be easily tuned by adjusting properties of the e-beam. Alanine is commonly used for radiation dosimetry, and the radicals generated upon irradiation are known to be stable for many years at room temperature if stored anhydrously. To improve thermal conductivity and permit efficient spin diffusion between polarising centres, irradiated polycrystalline alanine was dispersed in glycerol, resulting in radical lifetimes that were stable for several weeks when stored at room temperature with silica desiccant gel. Upon dissolution, radicals were quenched, yielding a non-toxic, directly injectable substrate. At 6.7 T, we were able to obtain nuclear polarisation levels of approximately 20% and discovered that the polarisation transfer is not simply explained by the solid effect or models of thermal mixing with the known stable alanine radical -factor and hyperfine couplings at this field strength. This is unlike similar analytic models of DNP with pyruvate and disperse chemical free radicals such as trityl[13]. Crucially, the hyperpolarised 13C label on alanine was successfully transferred to pyruvate and lactate after injection in vivo, demonstrating the potential of this method for metabolic imaging.
2 Results
2.1 Irradiation of Samples
Dry polycrystalline natural-abundance (NA) and [1-13C]L-alanine powders were irradiated using a FLASH-optimised[14] in-house developed 6 MeV (nominal energy) electron linear accelerator.[15] This specialised accelerator (Fig. 1a) was based on a reconfigured Elekta SL75-5 travelling-wave waveguide with an S-band RF magnetron source () and delivers ultra-high dose rate horizontal electron beam macropulses, with adjustable pulse repetition rates (25-) through precise control of RF frequency and injection parameters. [15] Unlike conventional medical linacs that employ bending magnets, our system delivered electrons in a straight-through beam path with specialised quadrupole focusing and advanced beam monitoring systems,[16, 17] allowing exquisite control of dose-per-pulse across multiple orders of magnitude, and was uniquely able to irradiate alanine samples, provided as a powder sample (Fig. 1b), with up to in seconds. Further details are provided in the SI §LABEL:SI:irradiation.
We undertook EPR spectroscopy to quantify radical species generated by electron irradiation. As radiation damage in dry, natural-abundance (NA) alanine is known to produce stable alanine radicals species (SARs) that have been extensively studied, both via EPR and theoretically,[18, 19, 20, 21, 22, 23, 24, 25] we compared experimental EPR spectra to that of TEMPO ((2,2,6,6-Tetramethylpiperidin-1-yl)oxyl), an aminoxyl radical used for DNP, and additionally quantitatively verified their dose/response curve. Endogeneous alanine radicals had a broader linewidth than TEMPO (Fig. 1c) and, as expected an increase in the irradiation dose led to a greater spin count, saturating at (Fig. 1d). We found that [1-13C]alanine had a drastically different EPR spectra compared to NA alanine, reflecting a large hyperfine interaction (not altering the corresponding spin count; SI §LABEL:SI:SpinCount). The -tensor and proton/nitrogen hyperfine interaction constants for NA alanine have previously been well documented;[26, 25, 22, 27] however, no such terms for [1-13C]alanine have been measured. We determined using a constrained Bayesian optimisation routine[28] the values of and for the predominant SAR species, known as R1 (Fig. 1f; SI §LABEL:SI:AlanineSAROptimisation); the principal values determined for were with Euler angles of in the previously reported molecular frame[22].
2.2 DNP via Irradiation-Induced Radicals
We first tested the feasibility of utilising the irradiation-induced radicals for DNP by exploring their polarisation characteristics and stability during sample storage. We investigated both dry powder mechanically secured in place with a separate snap-frozen “plug” of anhydrous glycerol and an alanine and glycerol mixture. We found that limiting nuclear polarisation increased with increasing radical concentration, but that significantly greater nuclear enhancement at was obtained with alanine crystallites dispersed in the glycerol matrix (Fig. 2a). These mixtures were stable for weeks independent of radiation dose, as quantified by EPR (Fig. 2b), with no significant change in concentration vs time. The addition of glycerol did not change the electronic environment at room temperature (Fig. 2c), but did at (Fig. 2d) with a shift in the electronic -tensor that we were not able to model.
Given the comparatively wide EPR linewidth of alanine radicals at , we elected to undertake DNP at . Under these conditions the relatively small degree of anisotropy would be expected to scale proportional to whereas the strong hyperfine couplings are field independent (depending primarily on the Fermi contact mechanism). We therefore expected to a first approximation an improvement in DNP efficiency through an analytic model of the thermal mixing process of polarisation growth based on the Provotorov equations,[29] as the effective width of the EPR lineshape would be expected to decrease. We additionally explored radiation doses (, , and ) to reflect the concentrations of broad line width exogenous chemical radical species typically used for DNP (i.e., 30 mM, 50 mM, and 70 mM). To our surprise, we did not observe the characteristic “bimodal” frequency sweep pattern associated with either the well-resolved solid effect or thermal mixing mechanisms, the latter of which is expected to describe DNP with nitroxide or trityl radicals and produces maxima approximately at the sum or difference of the nuclear and electron Larmor frequencies (). Instead we observed an asymmetric “hump” at higher frequency to the major lobe of the curve that decreased with increasing radiation dose (i.e. approximately at alone) as illustrated in Fig. 3a. At lower doses (), we observe a bimodal microwave profile with a prominent second lobe that progressively diminishes as the radiation dose increases to , and ultimately disappears entirely at . This systematic change suggests a transition in the underlying polarisation transfer mechanism that correlates directly with radical concentration. A biexponential build-up profile was also observed (Fig. 3b).
To understand this difference upon the addition of glycerol, we undertook X-ray diffraction. Data were obtained from natural abundance polycrystalline alanine and polycrystalline alanine dispersed in glycerol; a spherical harmonic Rietveld crystallographic texture refinement[30] aiming to quantify any degree of order in the sample was undertaken with MAUD.[31] Using a neutron-diffraction crystal structure obtained on the SXD instrument[32] at the ISIS neutron source,[33] we demonstrated strong partial ordering of the sample (Fig. 3c) that changed (and strengthened) upon addition of glycerol (Fig. 3d). As is common in texture analysis, we express orientation in units of multiples of a random distribution (mrd), where a sample without preferred orientation is one uniformly; alanine and glycerol mixtures had orientation preferences occurring and are partially ordered. This was consistent with ab initio liquid-state molecular dynamic simulations performed on two semi-infinite glycerol/alanine domains in LAAMPs[34] with EMC[35] that predicted the immiscibility of the two substances (SI §LABEL:SI:MolecularDynamics).
This (Fig. 3a) microwave sweep profile is distinct from others previously reported, and the mechanistic processes by which DNP occurs appear as efficient as trityl radicals used in clinical research. We therefore undertook control DNP experiments comprising substrates polarised via exogenous chemical radical trityl species, namely 13C-pyruvate with AH111501 and 13C-alanine with OX063 Trityl (Fig. 3e). We observed that the greater the radical concentration, the longer the slow component of the build-up time and the shorter the fast component (SI §LABEL:SI:builduptimes).
Given the wide EPR lineshape of the stable alanine radical, we investigated the use of a frequency-swept microwave pulse sequence for DNP. This resulted in an increased polarisation transfer and increased enhancement, within a factor of 2.5 of the “gold standard” pyruvate with trityl radical (Fig. 4a). In an attempt to understand the mechanism behind this observation, we simulated the fitted EPR lineshape under DNP conditions. We found that a polycrystalline distribution of R1 radicals have distinct spectral density functions depending on preferred orientation, quantified by the order parameter where the angle between the molecular -axis and is distributed by , with the orientation preference found by X-ray diffraction to correspond to a combination of these parameters (Fig. 4b). A detailed quantum-mechanical simulation of this pulse sequence together with the known radical parameters was then undertaken in the advanced magnetic resonance simulation environment Spinach.[36] Despite including more than 50 13C spins representing an alanine supercell with detailed measurements of its proton and carbon chemical shift tensors,[37, 26] simulating the entirety of the minute-long experiment at high temporal resolution (compared to the electron Larmor frequency) and utilising the Arcus-C supercomputing cluster,[38] we were unable to reproduce this experimental result and instead predicted a bimodal curve qualitatively similar to that expected for thermal mixing (SI §LABEL:si:spinach). Previously published analytic models of DNP via thermal mixing and the cross effect[29, 39, 40] likewise produce a predicted bimodal curve with either predicted analytic or experimental EPR data (SI §LABEL:si:analyticmodels). A pseudopotential plane-wave simulation in QuantumESPRESSO[41] with our structures predicts that alanine crystals are a direct-gap semiconductor with a large () bandgap (SI §LABEL:si:QE, experimentally validated[42]); were the effect of electron irradiation sufficient to dope spins into the conduction band, we would expect to see a well-resolved solid effect mechanism and narrow EPR data accordingly.
This, again, did not match the data. We hypothesise that the density of unpaired electrons is insufficient to form a delocalised conduction band but rather creates localised mid-gap states. This intermediate regime, where radicals are neither fully isolated molecules (as often assumed in molecular DFT calculations used for alanine radicals[22]) nor fully delocalised carriers (as in classical solid effect DNP systems such as paramagnetic salts[43]), may explain why neither approach fully captures our experimental observations. The emergence of a complex and strongly orientation-dependent EPR lineshape suggests that these localised states retain molecular character in diffuse partially ordered systems while experiencing significant inter-radical coupling mediated by the alanine crystal lattice and glycerol matrix, necessitating models that bridge molecular and solid-state treatments that we were unable to provide.
Nevertheless, encouraged by the feasibility of polarising [1-13C]alanine via irradiation-induced radicals, our focus turned to exploiting this for biomedical imaging. Empirically optimised DNP schemes compatible with our hardware produced an estimated solid state polarisation of approximately 20% (Fig. 4c), similar to that obtained with a separate sample of [1-13C]alanine and OX063 trityl radical, both of which are comparable to the “gold standard” used clinically of pyruvate with the trityl radical AH111501, which polarised to approximately 40%. After dissolution, the immediate quenching in aqueous media of endogenous alanine radicals resulted in an increase in the apparent liquid-state at (Fig. 4d), as may be expected from the removal of paramagnetic relaxation centres.[9]
2.3 In Vivo Application as a Metabolic Tracer
To test the optimised irradiated alanine preparation and DNP protocol, two proof of concept in vivo studies were conducted in rats. First, slice-selective magnetic resonance spectroscopy acquisitions were performed following separate tail-vein injections of [1-13C]alanine polarized with trityl radicals as described previously[44] or endogenous radicals generated via 70 kGy electron-beam irradiation. Temporal spectroscopic analysis demonstrated higher initial signal magnitude from the irradiated sample relative to the trityl-based preparation (Fig. 5a, b). Secondary resonances were observed flanking the primary [1-13C]alanine signal in both preparations. Spectral summation was performed to enhance the signal-to-noise ratio (SNR) and facilitate metabolite identification (Fig. 5c, d). The chemical shifts of these secondary resonances ( and ) corresponded precisely with the literature-reported values for [1-13C]lactate and [1-13C]pyruvate, respectively[45]. Subsequent metabolite quantification via the AMARES algorithm revealed comparable metabolic dynamics between both preparation methods (Fig. 5e, f) and spectral SNR was sufficient to quantify the downstream metabolites expected in vivo. No additional labelled peaks were observed; the products of the dissolution were identical to that observed with trityl radical.
Secondly, comparative 2D magnetic resonance spectroscopic imaging (MRSI) was performed in a separate animal using identical hyperpolarised [1-13C]alanine preparations. MRSI data were acquired, together with T2-weighted proton images acquired in the axial plane for anatomical localisation. Spatiotemporal analysis of the spectroscopic data demonstrated concordant biodistribution patterns between the trityl-based and irradiation-based preparations (Fig. 5g), with predominant accumulation in the renal parenchyma as alanine perfuses the kidneys, consistent with previous renal reports and enabling the technique to quantify redox potential,[44] and confirming equivalent in vivo behaviour between preparation methods.
3 Discussion
The clinical feasibility and utility of metabolic imaging with hyperpolarised MRI and dDNP has been consistently demonstrated since its inception 22 years ago [46], with the transformative ability to spatially map metabolism of huge interest in many disparate diseases including cancer,[1] heart disease,[47] and diabetes[48]. With more than 1000 human patients scanned and multi-centre trials soon to commence[49] it is a physical technique that has opened new avenues for research and the diagnosis and management of disease. However, for its integration within routine clinical workflows to be fully realised, significant improvements in its cost, accessibility and practicality are essential. Beyond the substantial capital expenditure for dDNP hardware (typically exceeding $2 million), institutions using dDNP face prohibitive ongoing operational costs. These include maintaining dedicated clean-room pharmacy facilities with specialised personnel for the sterile preparation of per-patient disposable fluid paths for dDNP, implementing expensive quality control measures, and adhering to stringent regulatory frameworks for pharmaceutical preparation. Each hyperpolarised examination currently requires the labour-intensive preparation of sterile fluid paths,[50] their filling with sterile stable-isotope labelled metabolites together with dissolution and neutralisation media (both manufactured according to sterile guidelines that vary regionally[49]) and a trityl radical that provides the unpaired electron the technique requires. The synthesis of trityl radicals is not trivial, typically requiring 13 separate organic synthetic steps, leading to a cost $10k/g.[6] Moreover, the required subsequent filtration steps to remove the radical add further complexity to the design of the fluid path. These are all individual processes that substantially increase per-patient costs and logistical complexity and restrict the availability of a technique that has the ability to directly measure central metabolic reactions key to all life. Furthermore, the limited shelf-life of the radical in prepared samples (approximately 24 hours at room temperature for trityl[51]) necessitates precise coordination between pharmacy operations and clinical scheduling, and prohibits the transport of dDNP samples nationally or internationally unless extraordinary steps are taken such as transporting samples cryogenically under defined (and varying) temperatures and magnetic fields.[52, 53, 12] As well as restraining human experiments, these large costs further constrain dDNP availability in non-human biological research: although of utility studying species such as snakes[45] and veterinary canine patients,[54] trityl radicals are sufficiently expensive as to be repurified from animal urine after use.[55]
To address these fundamental constraints on widespread adoption, we explored ultra-high dose rate electron-beam irradiation for generating inherently biologically sterile samples with radicals in situ: a sample could be prepared in a central (worldwide) facility, rapidly sterilised, and shipped at room temperature around the world prior to use with a polariser in a hospital site. Furthermore, the ability to use powders or crystals of neutral salts, such as alanine, rather than glassing suspensions of trityl radicals and typically organic ketoacids, further dramatically ameliorates the technical requirements of the technique. This approach also opens new classes of molecules to dDNP, as radical formation occurs statistically throughout the sample via well-understood particle-physics interactions, rather than requiring glassed mixtures of disparate radicals (themselves being limited by exclusion from any crystals formed). This allows for effective polarisation of high concentrations of virtually any anhydrous solid-state molecule that would otherwise be challenging to study. All that would be required for dissolution is a neutral sterile dissolution buffer, and, unlike UV-irradiation generated radicals in acids, there is no need for precise control of temperature and magnetic field over many orders of magnitude.[12] Indeed, in this work, we prepared samples for DNP in one country (the UK) and studied them via dissolution-DNP in another (Denmark).
As anticipated from extensive studies on NA alanine as a dosimeter, e-beam irradiation of NA alanine generated radicals in a dose-dependent manner[56], and a similar trend was observed from the 13C-enriched samples, saturating at around , well above the dose requirement of for medical sterilisation under ISO 11137-2.[57, 58] What is surprising, however, is that upon mixing with glycerol the system appears to become partially ordered, and the efficiency of DNP increases substantially. This also generates a clinically useful probe as alanine and glycerol are not miscible in the volume ratios considered here and stable alanine radicals remain present for at least 16 weeks, stored anhydrously in an inexpensive plastic container with silica desiccant gel at room temperature. Furthermore, changes at low temperature with this system, indicating a fundamental change in the radical electronic environment. We observed no unexpected radiation chemistry byproducts (likely because they spontaneously recombine upon the addition of hot water in the dissolution process, wherein the mixture is melted), and both glycerol and alanine are endogenous biomolecules – which are safe in vivo. This permitted the quantification in the rat kidney of downstream alanine metabolites, directly analogously to the reported biomedical uses of [1-13C]alanine hyperpolarised with trityl radicals.[59, 60, 44, 61, 62]
This polarisation enhancement is difficult to quantitatively understand. We have discovered a unique DNP mechanism that is not explained by conventional models yet produces clinically relevant levels of polarisation in an extensively studied biomolecule. The alteration of the frequency sweep profile of nuclear enhancement with radiation dose we hypothesise may be due to some degree of cooperativity between electrons. The temperature-dependent electronic -shift we observed in the EPR data raises intriguing questions about potential magnetic ordering in our system. Indeed, spontaneous spatial ordering in neutral alanine crystallites has previously been reported to occur in strong magnetic fields,[63] arising out of magnetic anisotropy of the L-alanine unit cell and its comparatively large (in magnitude) ;[64, 65, 66] and zirconium-nitrate doped -alanine single crystals form magnetic order and have been used as nonlinear optical devices.[42]
While traditional paramagnetic systems maintain independent spin behaviour even at low temperatures, our partially ordered alanine crystallites dispersed in glycerol could exhibit more complex magnetic phenomena as radical centres become close enough to develop short-range magnetic correlations at low temperatures and high radiation doses within individual crystallites. The glycerol matrix in this scenario could function analogously to domain boundaries in ferromagnetic materials. By physically separating the alanine crystallites, glycerol creates discontinuities in the exchange pathways, limiting the correlation length of any magnetic ordering, but permitting thermodynamically driven spin diffusion through weak coupling provided by proton/carbon spins in the matrix. This is conceptually similar to the role of Bloch walls in ferromagnetic materials, though operating at the nano- or microscale between distinct crystallite domains rather than within a continuous bulk material. This model could potentially explain several of our observations:
Firstly, the spatial confinement of magnetic correlations to individual crystallites would naturally lead to anisotropic magnetic responses dependent on crystallite orientation, consistent with our texture analysis showing preferential alignment. Each crystallite may develop its own local magnetic environment influenced by its specific orientation relative to the external field.
Secondly, this partial magnetic ordering could modify the effective field experienced by nuclear spins, altering the resonance conditions for polarisation transfer for that microscopic environment. In conventional DNP models, polarisation transfer is mediated by electron-nuclear dipolar interactions under well-defined frequency matching conditions (e.g. for the solid effect) and the angular orientation terms in the dipolar interaction Hamiltonian () are assumed to be spatially averaged in a glass. This regionality in orientation together with a transition from individual to collective behaviour would explain why the best conventional theoretical models of DNP fail to predict our observations at higher radical concentrations, as they assume only electron-nuclear interactions and/or are based on thermodynamic arguments assuming some degree of spatial homogeneity and good thermal contact. They do not account for many-body effects that emerge when multiple electrons (arising from multiple radicals in close proximity) interact collectively.
Thirdly, the glycerol “boundaries” could create a heterogeneous distribution of magnetic environments throughout the sample. Radicals near the glycerol interface might experience different effective fields than those in crystallite interiors, leading to a distribution of resonance conditions that broadens and shifts the optimal DNP frequency. This domain-like structure might also explain the effectiveness of frequency-swept microwave pulses in our system. By sweeping across a range of frequencies, we could be sequentially addressing different sub-populations of radicals with distinct magnetic environments, each contributing to the overall polarisation enhancement. The constructive summation of these contributions could explain the superior performance compared to fixed-frequency irradiation, even though the frequency swept range is small compared to the overall EPR lineshape. It also provides a mechanism for the biexponential growth of polarisation, which could be ascribed as being that within each crystallite occurring quickly, and that between them occurring through the dilute matrix of natural abundance carbon-13 spins in glycerol. These results are overall surprising, and future work will help elucidate a mechanism behind them. They represent a new paradigm in DNP, where a partial spatial ordering of spin systems creates emergent behaviour not predicted by isolated electron models of molecular physics.
The levels of nuclear enhancement we achieved are several orders of magnitude larger than those reported for ionising-radiation radical induced DNP previously by -rays by Giannoulis,[10] and are comparable with the optimal trityl, narrow line-width radicals used clinically. For a non-irradiated alanine sample, the polarisation level at the point of dissolution was 19.2% and the nuclear T1 was . For reference, a previous study where alanine was also hyperpolarised with the trityl radical yielded a liquid state polarisation of 12.6% at the point of dissolution and an in vitro T1 of at [60]. Our irradiated samples produced comparable results; polarisation levels of 17.7% at the point of dissolution were achieved, while the extracted nuclear T1 relaxation times were at . The nuclear T1 relaxation time was prolonged due to the fact that the irradiation-generated radicals quench upon dissolution. Future work could explore the direct deuteration of the 13C-enriched alanine to potentially enhance polarisation levels as previously observed for other molecules hyperpolarised by dDNP[67]. Crucially, the pH of the irradiated alanine samples (approximately 7) also made them suitable for in vivo trial with dissolution in water; we note that neutral salts are readily crystallised and substances such as sodium pyruvate likely will be readily amenable to this technique
Beyond the physical novelty of our approach, our in vivo experiments demonstrate significant translational potential. Electron-irradiated 13C-alanine not only matched but in some aspects exceeded the performance of conventional preparations using exogenous radicals. It is worth considering how 13C-alanine could compare to 13C-pyruvate, the most commonly used metabolic probe with dDNP and undergoing clinical trials. Whilst pyruvate polarises very well with dDNP and is rapidly metabolised to lactate, alanine, and bicarbonate, it has been noted that the pyruvate/lactate ratio is not reflective of intracellular redox state following HP pyruvate infusion[68]. Upon switching the injectable substrate from 13C-pyruvate to 13C-alanine, Hu et al. recorded more than a ten-fold increase in the 13C-lactate/13C-pyruvate ratio[60]. The difference was suggested to be due to the uptake of 13C-pyruvate being limited by perfusion and downstream flux through monocarboxylate transporters during the acquisition window. An incomplete uptake of 13C-pyruvate would mean contributions from an extracellular 13C-pyruvate pool in the measured 13C-pyruvate signal. Furthermore, any extracellular 13C-pyruvate present in plasma can be converted into 13C-lactate by red blood cells that contain lactate dehydrogenase, leading to an “extracellular” 13C-lactate pool[59]. Signals arising from intracellular and extracellular pools cannot be easily distinguished in hyperpolarised MR, leading to the proposed use of 13C-alanine for determining the 13C-lactate/13C-pyruvate ratio. For 13C-alanine to be converted into 13C-pyruvate, and subsequently 13C-lactate, it must first be exposed to the intracellular enzyme alanine transaminase. Thus, the 13C-pyruvate signal detected following the injection of 13C-alanine would be intracellular in origin. Moreover, after the metabolism of alanine is observed and quantified, the cellular redox potential can be easily determined [59] from the known relation where .
Our electron irradiation approach uniquely combines this biochemical advantage with the practical benefits of in situ radical generation and spontaneous quenching, providing a powerful new tool for quantifying intracellular redox potential. It is also the first to explicitly consider partially ordered solids as substrates for biomedical dDNP rather than amorphous glasses that are commonly used with the technique (and, indeed, often listed as a requirement for its success; [69]) While this technique could be extended to other substrates, 13C-alanine represents an ideal initial candidate due to its metabolic significance, detailed quantum-mechanical study, and demonstrated biochemical and clinical relevance. This approach may reduce the technological, regulatory, and economic barriers that have constrained hyperpolarised MRI primarily to specialised research centres, potentially expanding access to this powerful diagnostic capability.
3.1 Limitations
We should note that several limitations should be acknowledged in this work: notably, our mechanistic understanding of the DNP process in this system remains incomplete and in need of extension via theoretical modelling. The unusual frequency sweep profiles of this highly efficient polarisation process hint at cooperativity and physics beyond conventional DNP models, but a comprehensive theoretical framework that quantitatively predicts these behaviors is lacking. Additional spectroscopic studies with variable temperature, field strength, and radical concentration would help elucidate the underlying mechanisms. Additionally, while we demonstrated stable radical formation in alanine, the applicability of this approach to other metabolically relevant molecules requires further investigation. Different molecular structures may respond differently to electron irradiation, potentially generating distinct radical species with varying stability and DNP efficiency. The specific arrangement of alanine in a crystalline lattice likely contributes to the observed radical stability and its partial ordering in glycerol (analogous to a nemantic liquid crystal), and molecules with different crystal structures, or those that do not readily crystallise, may present challenges. Furthermore, the comparatively high radiation doses considered may embrittle plastics used in the construction of fluid paths, but this phenomenon is well-characterised, and many medical plastics are routinely sterilised with comparable doses.[70, 71]
Finally, the highest nuclear polarisation levels achieved with our approach () remain lower than those reported for optimised pyruvic acid preparations with trityl radicals (). While sufficient for in vivo imaging, higher polarisation would improve SNR and potentially allow detection of lower-concentration metabolites. Further optimisation of irradiation parameters, alanine crystallite size, and DNP conditions might narrow this gap. Likewise, owing in part to this limitation, our in vivo experiments were limited to proof-of-concept demonstrations in a small number of healthy animals. The performance of irradiated alanine in disease models, particularly its sensitivity to pathological alterations in redox potential, requires validation in relevant preclinical models before clinical translation.
4 Methods
4.1 Sample Preparation
The samples used in this study can be categorised into three main groups: alanine containing endogenous irradiation-generated radicals, alanine mixed with the exogenous trityl radical OX063 at a concentration of , and pyruvate with AH1111501. These exogeneous radicals differ only by an exterior methyl ‘R’-group in their chemical structure and have almost identical electronic properties; these concentrations of both trityl radicals were chosen to reflect reported optimal conditions for DNP for both molecules.
To prepare the former, dry [1-13C]L-alanine powder (Isotec and Cambridge Isotope Laboratories, Inc) was irradiated using an ultra-high dose rate 6 MeV electron linear accelerator operating in pulsed mode, delivering of dose to the sample in each of the pulses, with a pulse repetition rate of and an approximate beam current of . The dose of irradiation was varied between 10 – depending on the desired radical concentration and was linear with the total number of pulses delivered. The dosimetry and beam delivery was controlled using a beam monitor, which has been described in detail elsewhere,[16, 17] and calibrated against radiochromic film (EBT-XD, Ashland Inc., Covington, KY, USA). The total time for irradiation was deliberately lengthened to avoid linac heating; approximately was irradiated in . The irradiated powder was then used as the main component for two different sample types: either an irradiated dry powder or an irradiated powder/anhydrous glycerol mix with a 1:1 w/w ratio.
The alanine preparation containing OX063 trityl radical (molecular weight = ; GE Healthcare) comprised of [1-13C]L-alanine mixed in 8 ml of water and of HCl. This mixture was freeze-dried, and approximately of alanine hydrochloride yielded. The alanine hydrochloride was then dissolved in DMSO with the assistance of heating. OX063 trityl was added to achieve approximately radical concentration. of this mixture was pipetted into a sample cup.
For the pyruvate sample, [1-13C]pyruvic acid was mixed with the appropriate mass of AH111501 trityl (GE Healthcare; CID 11607875) to achieve a radical concentration, of which was pipetted into a sample cup.
4.2 Dynamic Nuclear Polarisation
Samples were polarised at 3.35 T or 6.7 T using a HyperSense (Oxford Instruments) or SpinAligner (Polarize), respectively. Further details are provided in the SI §LABEL:si:DNP.
4.3 In Vitro Experiments
The polarisation levels and nuclear T1 relaxation times of samples were recorded using a benchtop NMR spectrometer at 1.4 T (SpinSolve 60ULTRA Carbon, Magritek). The accompanying software also enabled the level of polarisation to be estimated at the point of dissolution.
4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall.
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, slice thickness was used to ensure the whole abdomen of the rat was imaged. of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A flip angle was applied with a repetition time (TR) of for a total of 5 minutes (100 time points). For the MRSI, of hyperpolarised substrate was injected over approximately , and the scan started upon the start of the injection. A matrix was chosen, and a FA was applied with a TR of for a total of 2 minutes (8 time points and approximately per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin) was used to record EPR spectra at room temperature. For recording spectra at and an estimate of , an X-band CW spectrometer (E680, Bruker BioSpin) was used.
Magnesium Oxide powder, a common standard reference in EPR,[72] was used to account for changes in the magnetic field between experiments due to differences in sample positioning.
Samples that comprised dry powder were directly loaded into thin wall quartz EPR tubes, and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25–. Liquid state samples were first loaded in diameter capillary tubes, which in turn were loaded into the EPR tubes.
Spin counts were determined from EPR spectra using an in-house developed MATLAB script. The script incorporated baseline correction, normalised for differences in resonator -factor, and integrated the recorded spectra.
4.6 X-ray Diffraction
XRD was conducted using copper X-rays, of wavelength and , in an Empyrean (Malvern Panalytical) diffractometer. XRD samples comprised dry alanine powder or an alanine powder/glycerol mix (1:1 w/w). The Cu wavelength was removed from the diffraction signal with a Ni filter placed in front of the PIXcel1D detector. A set of programmable divergence slits were used for the incident optic, set to illuminate a fixed surface area of across the angular range of the scan. During the scans, the samples were rotated about the azimuthal axis at a frequency of to fully characterise the crystallite distribution.
5 Conflict of Interest
JYCL, CL, JJM, CR and DT are listed as authors on a filed patent (EP24168322.6) pertaining to the use of ultra-high dose rate electron irradiation for preparing hyperpolarised contrast agents.
Acknowledgments
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to J.J.M. and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the Department of Physics Mechanical Workshops is also kindly noted.
Author Contributions
C.H.E.R. and J.Y.C.L. conceptualised research, conducted experiments, analysed data, and drafted the manuscript. E.S.S.H., N.V.C., and D.A.D. assisted with sample preparation and in vivo/high field DNP experiments; B.W.C and S.S. assisted with low-field DNP experiments. K.P., I.T., B.V., C.H.E.R. undertook electron linear accelerator experiments for sample irradiation, including design and manufacture of equipment. J.L, C.H.E.R. and J.J.M undertook X-ray diffraction experiments and their interpretation. W.M., and A.B. provided expertise in EPR spectroscopy and assisted with data interpretation. Z.R. assisted with modelling. L.B.B. and C.L. supervised the in vivo experiments and contributed to experimental design. J.Y.C.L, D.J.T. and J.J.M. conceived and supervised the project, secured funding, analysed data, and finalized the manuscript. All authors reviewed, edited and approved the final manuscript.
Data Availability Statement
Raw CIF files for alanine in this phase are enclosed as SI with this paper. All other data is available from the authors upon reasonable request; request should be directed at the corresponding author.