### v2

# Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation Short Title: Electron-irradiated radicals for DNP

Abstract
Dissolution Dynamic Nuclear Polarisation (dDNP) increases the sensitivity of magnetic resonance experiments by $10^{4}-$fold, permitting isotopically-labelled molecules to be transiently visible in MRI scans. dDNP mechanistically takes place at $\sim 1$ K, and requires unpaired electrons and microwaves. These electrons are usually chemical radicals, requiring removal by filtration prior to injection into humans. Alternative sources, such as UV irradiation, generate lower polarisation and require cryogenic transport. We present ultra-high-dose-rate electron irradiation as a novel alternative for generating non-persistent radicals in alanine/glycerol mixtures. These are stable for months at room temperature, quench spontaneously upon dissolution, are present in dose-dependent concentrations, and generate comparable nuclear polarisation (17%) to trityl radicals used clinically (19%) through a novel mechanism we believe to involve partial ordering, and electron-electron interactions. Owing to the large radiation doses required, this process is sterilising, permits imaging of alanine metabolism in vivo in the rat kidney, and may aid clinically translating dDNP.
Teaser:
High-energy electrons generate stable radicals for safer, more accessible metabolic imaging with dDNP through a novel mechanism.
Main text
1 Introduction
Hyperpolarised Magnetic Resonance Imaging (HP MRI) is a molecular imaging technique that is predominantly used to monitor metabolic fluxes in real-time by transiently making a labelled molecular probe visible to magnetic resonance experiments such as MRI. Over the past two decades, its pre-clinical use has spanned multiple fields of basic scientific and medical research, with 2013 marking the start of its validation for various clinical applications, including prostate cancer (?) (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), traumatic brain injury (?) (?, ?, ?), heart failure (?, ?, ?, ?, ?, ?, ?, ?) and ischaemic heart disease, (?, ?, ?, ?) amongst many others. (?) To perform HP MRI, a hyperpolariser is required – a device that transiently enhances the MR signal of stable labelled isotopes by several orders of magnitude. One method to achieve this is through dissolution Dynamic Nuclear Polarisation (dDNP), which employs cryogenic temperatures, high magnetic field strengths (>3>3 T), and microwave irradiation. A typical sample for dDNP comprises a metabolic substrate mixed with a chemical radical species and optionally also a glassing agent such as glycerol. At low ($\sim 1$ K) temperature, the unpaired electron spins present in the radical become thermally polarised. When irradiated with microwaves, a population inversion of these electronic spins can be created, and they are driven to relax, primarily through dipolar interactions, and ultimately transfer their polarisation to nearby nuclei. The net process behind DNP therefore is that nuclear polarisation above that at thermal equilibrium is created. Once the desired nuclear signal enhancement is achieved, a heated and pressurised solvent is introduced to rapidly melt the solid-state sample and convert it into a form suitable for injection into a living system. Following injection, the enzymatic conversion of the hyperpolarised substrate into other metabolic intermediates can be temporally and spatially monitored in vivo through magnetic resonance experiments such as MRI.
[…]
However, regardless of whether the electron source is chemical or photon-based, there are substantial technical limitations on the ability of clinical sites to ship and receive compounds for hyperpolarisation with the radicals in situ. For instance, the chemical instability of the trityl radical — the most commonly used chemical radical — renders centralised production of pre-filled fluid paths for global distribution infeasible; fluid paths are transported typically at $195$  or 253 K. (?, ?, ?) In contrast, while low-energy gamma-generated radicals are reported to be stable at room temperature for several months to produce a single sample, irradiation can take days to months, and reported nuclear enhancement factors are on the order of $10^{2}$ and not the $10^{5}$ obtained with chemical radicals. (?) DNP with UV-generated radicals is only possible if samples are optically transparent (e.g. pyruvic acid) and require low-temperature (cryogenic) transportation (?), which incurs a substantial expense and is not routinely performed as precise control of magnetic field $B$ and temperature $T$ is required over many orders of magnitude. (?) Ultimately, sites conducting human dDNP experiments generally require local access to a pharmacy equipped with sterile compounding facilities to comply with the high degree of sterility assurance required. This is complex from a regulatory view, (?) and is considered a substantial financial burden that may ultimately limit the accessibility of the technique.
Here, we present an alternative approach for generating sterilised samples for clinical HP MRI that are suitable for convenient transport and centralised manufacture, potentially offering an alternative for sites that do not have access to a pharmacy. Using an ultra-high dose rate 6 MeV electron linear accelerator (linac), we demonstrate the creation of endogenous radicals within polycrystalline alanine, the concentration of which can be easily tuned by adjusting properties of the e-beam. Alanine is commonly used for radiation dosimetry, and the radicals generated upon irradiation are known to be stable for many years at room temperature if stored anhydrously. (?, ?) To improve thermal conductivity and permit efficient spin diffusion between polarising centres, irradiated polycrystalline alanine was dispersed in glycerol, resulting in radical lifetimes that were stable for several weeks when stored at room temperature with silica desiccant gel. Upon dissolution, radicals were quenched, yielding a non-toxic, (?, ?) directly injectable substrate. At 6.7 T, we were able to obtain nuclear polarisation levels of approximately 20% and discovered that the polarisation transfer is not simply explained by the solid effect or models of thermal mixing with the known stable alanine radical (SAR) $g$-factor and hyperfine couplings at this field strength. This is unlike similar analytic models of DNP with pyruvate and disperse chemical free radicals such as trityl (?). Crucially, the hyperpolarised 13C label on alanine was successfully transferred to pyruvate and lactate after injection in vivo, demonstrating the potential of this method for metabolic imaging.
2 Results
2.1 Irradiation of Samples
[…]
[Figure 1 approximately here]
We undertook Electron Paramagnetic Resonance (EPR) spectroscopy to quantify radical species generated by electron irradiation. As radiation damage in dry, natural-abundance (NA) alanine is known to produce stable alanine radicals species (SARs) that have been extensively studied, both via EPR and theoretically, (?, ?, ?, ?, ?, ?, ?, ?) we compared experimental EPR spectra to that of TEMPO ((2,2,6,6-Tetramethylpiperidin-1-yl)oxyl), an aminoxyl radical used for DNP, and additionally quantitatively verified their dose/response curve. Endogeneous alanine radicals had a broader linewidth than TEMPO (Fig. 1C) and, as expected an increase in the irradiation dose led to a greater spin count, saturating at $\sim$70 kGy (Fig. 1D) for the irradiated sample. We found that [1-13C]alanine had a drastically different EPR spectra compared to NA alanine, reflecting a hyperfine interaction to the 13C atom (not altering the corresponding spin count; SI §6.7). The $g$-tensor and proton/nitrogen hyperfine interaction constants $A$ for NA alanine have previously been well documented. (?, ?, ?, ?) however, no anisotropic such terms for [1-13C]alanine have been measured experimentally, although they have been predicted by Density Functional Theory calculations. (?) The EPR spectra obtained for both the NA and labelled samples at room temperature can be modelled reasonably using published hyperfine coupling values and a weighted combination of two SARs (cf. SI §6.7). For the purposes of predicting DNP behaviour, a single dominant species was considered and we modelled our system as a powder containing the labelled predominant SAR species, known as R1 (13C-SAR1). We determined using a constrained Bayesian optimisation routine (?) the values of the hyperfine coupling tensor to the 13C, $A_{13C}$, and the $g$-tensor(Fig. 1F; SI §6.8); the principal values determined for $A_{13C}$ were [1.80266,15.8427, 64.6933][-1.80266,\,-15.8427,\,64.6933] MHz with Euler angles of [1.50914, 1.30749, 0.761318][1.50914,\,1.30749,\,0.761318] rad in the previously reported molecular frame, (?) and these are a similar order of magnitude to the previously reported theoretical values calculated via Density Functional Theory (DFT), (?) although we note the accuracy of the $A_{13C}$ hyperfine values obtained is limited by the resolution of the linewidth of the EPR spectrum obtained. We additionally measured the electronic $T_{1e}$ and apparent $T_{2}$, $T_{m}$ at 5 K, and found that 100 kGy irradiated alanine had a long $T_{1}$ (7.6 s) in comparison to OX063 (1.4 s) but similar $T_{m}$ (332 ns vs. 370 ns; c.f. SI §6.7).
2.2 DNP via Irradiation-Induced Radicals
We first tested the feasibility of utilising the irradiation-induced radicals for DNP by exploring their polarisation characteristics and stability during sample storage. We investigated both dry powder mechanically secured in place with a separate snap-frozen “plug” of anhydrous glycerol and an alanine and glycerol mixture. This plug prevented the dry alanine powder from being displaced by any turbulent gas flow present, and additionally matched the dielectric loading conditions of the alanine/glycerol mixture, preserving RF coil loading and microwave power delivery between the two conditions. We further explored the comparatively limited effect of altering microwave power; c.f. SI §6.3. We found that the obtained nuclear polarisation increased with increasing radical concentration, and that substantially greater nuclear enhancement at 3.35 T was obtained with alanine crystallites dispersed in the glycerol matrix (Fig. 2A). These mixtures were stable for weeks independent of radiation dose, as quantified by EPR (Fig. 2B), with no significant change in concentration vs time (c.f. SI §6.4). The addition of glycerol did not change the electronic environment at room temperature as seen by EPR (Fig. 2C), but did at 5 K (Fig. 2D) with substantially altered hyperfine interaction constants and a shift in the mean isotropic electronic $g$-value (from $g=2.003698$ to $g=2.002862$, shown integrated in Fig. S11) that we were not able to model. The shape of the EPR spectrum at 5 K is likely indicative of a larger hyperfine interaction to the 13C atom, similar to that previously observed for the primary radical species which is only stable at low temperature; (?) indicating a further distribution over interconverting molecular species (c.f. SI §6.8). A frequency sweep at 3.35 T (Figure. S2) was difficult to interpret, with a small lobe of nuclear enhancement potentially present far away from the central maximum.
[Figure 2 approximately here]
Given the comparatively wide EPR linewidth of alanine radicals at 3.35 T, we elected to undertake DNP at 6.7 T. Under these conditions the relatively small degree of $g$ anisotropy would be expected to scale proportional to $B_{0}$ whereas the strong hyperfine couplings are field independent (depending primarily on the Fermi contact mechanism). We therefore expected to a first approximation an improvement in DNP efficiency through an analytic model of the thermal mixing process of polarisation growth based on the Provotorov equations, (?) as the effective width of the EPR lineshape $D$ would be expected to decrease, and the resolvable resolution of the DNP frequency sweep serve would, we believe, be predicted to increase at higher field. We additionally explored radiation doses (30 kGy, 50 kGy, and 70 kGy) to reflect the concentrations of broad line width exogenous chemical radical species typically used for DNP (i.e., 30 mM, 50 mM, and 70 mM). To our surprise, we did not observe the characteristic “bimodal” frequency sweep pattern associated with either the well-resolved solid effect or thermal mixing mechanisms, the latter of which is expected to describe DNP with nitroxide (?) or trityl (?) radicals and produces maxima approximately at the sum or difference of the nuclear and electron Larmor frequencies ($\omega_{e}\pm\omega_{N}$). Instead we clearly observed an asymmetric “hump” at higher frequency to the major lobe of the curve that decreased with increasing radiation dose (i.e. approximately at $\omega_{e}$ alone) as illustrated in Fig. 3A. At lower doses (30 kGy), we observe a bimodal microwave profile with a prominent second lobe that progressively diminishes as the radiation dose increases to 50 kGy, and ultimately disappears entirely at 70 kGy. This systematic change suggests a transition in the underlying polarisation transfer mechanism that correlates directly with radical concentration. A biexponential build-up profile was also observed (Fig. 3B) following irradiation at the maximum microwave frequency, with a fast component decreasing from 22.9 to 14.9 s and a slow component increasing from 0.92 to 6.88 hours as the radiation dose increased from 30 to 70 kGy (c.f. Table S2). This trend correlates with the disappearance of the bimodal frequency profile: samples exhibiting pronounced bimodal behavior (30 kGy) show faster overall polarisation kinetics, while those with unimodal profiles (70 kGy) exhibited slower kinetics. Whilst some degree of biexponential behaviour is expected for dilute electronic sources owing to spin diffusion, (?, ?) the longer time constant would decrease with increasing radical concentration – which we did not observe. A summary of fitted time constant parameters is provided in Table 1.
To understand this difference upon the addition of glycerol, we undertook X-ray diffraction. Data were obtained from natural abundance polycrystalline alanine and polycrystalline alanine dispersed in glycerol; a spherical harmonic Rietveld crystallographic texture refinement (?) aiming to quantify any degree of order in the sample was undertaken with MAUD. (?) Using a neutron-diffraction crystal structure obtained on the SXD instrument (?) at the ISIS neutron source, (?) we demonstrated strong partial ordering of the sample (Fig. 3C) that changed (and strengthened) upon addition of glycerol (Fig. 3D). As is common in texture analysis, we express orientation in units of multiples of a random distribution (mrd), where a sample without preferred orientation is one uniformly; alanine and glycerol mixtures had orientation preferences occurring >10>10 mrd and are partially ordered. This was consistent with ab initio liquid-state molecular dynamic simulations performed on two semi-infinite glycerol/alanine domains in LAAMPs (?) with EMC (?) that predicted the immiscibility of the two substances (SI §6.10).
[Figure 3 approximately here]
This (Fig. 3A) microwave sweep profile is distinct from others previously reported, and the mechanistic processes by which DNP occurs appear as efficient as trityl radicals used in clinical research. We therefore undertook control DNP experiments comprising substrates polarised via exogenous chemical radical trityl species, namely 13C-pyruvate with 30 mM AH111501 and 13C-alanine with 38 mM OX063 Trityl (Fig. 3E). We observed that the greater the radical concentration, the longer the slow component of the build-up time and the shorter the fast component (SI §6.5).
Given the wide EPR lineshape of the stable alanine radical, we investigated the use of a frequency-swept microwave pulse sequence for DNP. This resulted in an increased polarisation transfer and increased enhancement, within a factor of 2.5 of the “gold standard” pyruvate with trityl radical (Fig. 4A). In an attempt to understand the mechanism behind this observation, we simulated the fitted EPR lineshape under DNP conditions by modelling our room temperature data as containing 13C-SAR-1. We found that a polycrystalline distribution of R1 radicals have distinct spectral density functions $g(\omega)$ depending on preferred orientation, quantified by the order parameter $\lambda$ where the angle $\alpha$ between the molecular $z$-axis and $B_{0}$ is distributed by $p(\alpha)=\exp\left(-\lambda\left(3\cos^{2}\alpha-1\right)/2\right)$, with the orientation preference found by X-ray diffraction to correspond to a combination of these parameters (Fig. 4B), whose sum could approximately trace out the enhancement curve obtained. A detailed quantum-mechanical simulation of this pulse sequence together with the known radical parameters was then undertaken in the advanced magnetic resonance simulation environment Spinach. (?) Despite including more than 50 13C spins representing an alanine supercell with detailed measurements of its proton and carbon chemical shift tensors, (?, ?) simulating the entirety of the minute-long experiment at high temporal resolution (compared to the electron Larmor frequency) and utilising the Arcus-C supercomputing cluster, (?) we were unable to reproduce this experimental result and instead predicted a bimodal curve qualitatively similar to that expected for thermal mixing (SI §6.11). Previously published analytic models of DNP via thermal mixing and the cross effect (?, ?, ?) likewise produce a predicted bimodal curve with either predicted analytic (from our model) or experimental EPR data (SI §6.12). A pseudopotential plane-wave simulation in QuantumESPRESSO (?) with our structures predicts that alanine crystals are a direct-gap semiconductor with a large (5 eV) bandgap (SI §6.13, experimentally validated (?)); were the effect of electron irradiation sufficient to dope spins into the conduction band, we would expect to see a well-resolved solid effect mechanism. This, again, did not match the data. We hypothesise that the density of unpaired electrons is insufficient to form a delocalised conduction band but rather creates localised mid-gap states. This intermediate regime, where radicals are neither fully isolated molecules (as often assumed in molecular DFT calculations used for alanine radicals (?)) nor fully delocalised carriers (as in classical solid effect DNP systems such as paramagnetic salts (?)), may explain why neither approach fully captures our experimental observations. Another possibility is the competition of multiple, competing radical species that interact and interconvert, providing a superposition of radical environments. Further work is required to assess exactly the degree remaining of molecular character in these complex, diffuse, partially ordered systems which experience significant inter-radical coupling mediated by the alanine crystal lattice and glycerol matrix, necessitating models that bridge molecular and solid-state treatments that we were unable to provide.
Nevertheless, encouraged by the feasibility of polarising [1-13C]alanine via irradiation-induced radicals, our focus turned to exploiting this for biomedical imaging. Empirically optimised DNP schemes compatible with our hardware produced an estimated solid state polarisation of approximately 20% (Fig. 4C), similar to that obtained with a separate sample of [1-13C]alanine and OX063 trityl radical, both of which are comparable to the “gold standard” used clinically of pyruvate with the trityl radical AH111501, which polarised to approximately 40% (c.f. C). After dissolution, the immediate quenching in aqueous media of endogenous alanine radicals (?) resulted in an increase in the apparent liquid-state $T_{1}$ at 1.4 T (Fig. 4D), as may be expected from the removal of paramagnetic relaxation centres. (?)
[Figure 4 approximately here]
2.3 In Vivo Application as a Metabolic Tracer
To test the optimised irradiated alanine preparation and DNP protocol, two proof of concept in vivo studies were conducted in rats. First, slice-selective magnetic resonance spectroscopy acquisitions were performed following separate tail-vein injections of [1-13C]alanine polarised with trityl radicals as described previously (?) or endogenous radicals generated via 70 kGy electron-beam irradiation. Temporal spectroscopic analysis demonstrated higher initial signal magnitude from the irradiated sample relative to the trityl-based preparation (Fig. 5A, B). Secondary resonances were observed flanking the primary [1-13C]alanine signal in both preparations. Spectral summation was performed to enhance the signal-to-noise ratio (SNR) and facilitate metabolite identification (Fig. 5C, D). The chemical shifts of these secondary resonances (183 ppm and 171 ppm) corresponded precisely with the literature-reported values for [1-13C]lactate and [1-13C]pyruvate, respectively (?, ?, ?, ?). Subsequent metabolite quantification via the AMARES algorithm revealed comparable metabolic dynamics between both preparation methods (Fig. 5E, F) and spectral SNR was sufficient to quantify the downstream metabolites expected in vivo, although the comparatively long TR of 3 s precluded a detailed study of downstream kinetics whilst preserving magnetisation. No additional labelled peaks were observed; the products of the dissolution were identical to that observed with trityl radical (c.f. SI §6.14). The dissolution was ‘clean’, meaning that the entire volume of the sample was dissolved by the 5 ml D2O used as a dissolution buffer with no remaining partially dissolved sample remaining, and a final concentration of approximately 100 mM dissolved alanine was obtained post-dissolution, and subsequently 1 ml was manually injected over 5 s.
[Figure 5 approximately here]
Secondly, comparative 2D magnetic resonance spectroscopic imaging (MRSI) was performed in a separate animal using identical hyperpolarised [1-13C]alanine preparations. MRSI data were acquired, together with $T_{2}$-weighted proton images acquired in the axial plane for anatomical localisation. Spatiotemporal analysis of the spectroscopic data demonstrated concordant biodistribution patterns between the trityl-based and irradiation-based preparations (Fig. 5G), with predominant accumulation in the renal parenchyma as alanine perfuses the kidneys, consistent with previous renal reports and enabling the technique to quantify redox potential, (?) and confirming equivalent in vivo behaviour between preparation methods.
3 Discussion
The clinical feasibility and utility of metabolic imaging with hyperpolarised MRI and dDNP has been consistently demonstrated since its inception 22 years ago  (?), with the transformative ability to spatially map metabolism of huge interest in many disparate diseases including cancer, (?) heart disease, (?) and diabetes (?). With more than 1000 human patients scanned and multi-centre trials soon to commence (?) it is a physical technique that has opened new avenues for research and the diagnosis and management of disease. However, for its integration within routine clinical workflows to be fully realised, significant improvements in its cost, accessibility and practicality are essential. Beyond the substantial capital expenditure for dDNP hardware (typically exceeding \$2 million), institutions using dDNP face prohibitive ongoing operational costs. These include maintaining dedicated clean-room pharmacy facilities with specialised personnel for the sterile preparation of per-patient disposable fluid paths for dDNP, implementing expensive quality control measures, and adhering to stringent regulatory frameworks for pharmaceutical preparation. Each hyperpolarised examination currently requires the labour-intensive preparation of sterile fluid paths, (?) their filling with sterile stable-isotope labelled metabolites together with dissolution and neutralisation media (both manufactured according to sterile guidelines that vary regionally (?)) and a trityl radical that provides the unpaired electron the technique requires. The synthesis of trityl radicals is not trivial, typically requiring $\sim$13 separate organic synthetic steps, leading to a cost >>\$10k/g. (?) Moreover, the required subsequent filtration steps to remove the radical add further complexity to the design of the fluid path. The trityl radical used in human studies, AH111501, also known as EPA, is the methylated form of OX063, and is specifically engineered to precipitate out in acidic aqueous solution for filtration within the sterile fluid path. Indeed, to satisfy regulatory considerations it is currently a release criterion for all sites performing human hyperpolarised experiments that the radical concentration is below an agreed number, typically less than 3 mM in the dissolved liquid. (?) These are all individual processes that substantially increase per-patient costs and logistical complexity and restrict the availability of a technique that has the ability to directly measure central metabolic reactions key to all life.
Furthermore, the limited shelf-life of the radical in prepared samples (approximately 24 hours at room temperature for trityl (?)) necessitates precise coordination between pharmacy operations and clinical scheduling, (?) and prohibits the transport of dDNP samples nationally or internationally unless extraordinary steps are taken such as transporting samples cryogenically under defined (and varying) temperatures and magnetic fields. (?, ?, ?) As well as restraining human experiments, these large costs further constrain dDNP availability in non-human biological research: although of utility studying species such as snakes (?) and veterinary canine patients, (?) trityl radicals are sufficiently expensive as to be repurified from animal urine after use. (?)
To address these fundamental constraints on widespread adoption, we explored ultra-high dose rate electron-beam irradiation for generating inherently biologically sterile samples with radicals in situ: a sample could be prepared in a central (worldwide) facility, rapidly sterilised, and shipped at room temperature around the world prior to use with a polariser in a hospital site. Furthermore, the ability to use powders or crystals of neutral salts, such as alanine, rather than glassing suspensions of trityl radicals and typically organic ketoacids, further dramatically ameliorates the technical requirements of the technique, as a dissolution with sterile water is easier to obtain than with a carefully measured equimolar quantity of neutralising (basic) sterile buffer. This approach also opens new classes of molecules to dDNP, as radical formation occurs statistically throughout the sample via well-understood particle-physics interactions, rather than requiring glassed mixtures of disparate radicals (themselves being limited by exclusion from any crystals formed). This allows for effective polarisation of high concentrations of virtually any anhydrous solid-state molecule that would otherwise be challenging to study. All that would be required for dissolution is a neutral sterile dissolution buffer, and, unlike UV-irradiation generated radicals in acids, there is no need for precise control of temperature and magnetic field over many orders of magnitude. (?) Indeed, in this work, we prepared samples for DNP in one country (the UK) and studied them via dissolution-DNP in another (Denmark).
As anticipated from extensive studies on NA alanine as a dosimeter, e-beam irradiation of NA alanine generated radicals in a dose-dependent manner (?), and a similar trend was observed from the 13C-enriched samples, saturating at around 70 kGy, well above the dose requirement of 25 kGy for medical sterilisation under ISO 11137-2. (?, ?) What is surprising, however, is that upon mixing with glycerol the system appears to become partially ordered, and the efficiency of DNP increases substantially. This also generates a clinically useful probe as alanine and glycerol are not miscible in the volume ratios considered here and stable alanine radicals remain present for at least 16 weeks, stored anhydrously in an inexpensive plastic container with silica desiccant gel at room temperature. Furthermore, $g$ changes at low temperature with this system, indicating a fundamental change in the radical electronic environment. We observed no unexpected radiation chemistry byproducts (likely because they spontaneously recombine upon the addition of hot water in the dissolution process, wherein the mixture is melted), and both glycerol and alanine are endogenous biomolecules – which are safe in vivo. This permitted the quantification in the rat kidney of downstream alanine metabolites, directly analogously to the reported biomedical uses of [1-13C]alanine hyperpolarised with trityl radicals. (?, ?, ?, ?, ?)
This polarisation enhancement is difficult to quantitatively understand. We have discovered a unique DNP mechanism that is not explained by conventional models yet produces clinically relevant levels of polarisation in an extensively studied biomolecule. The alteration of the frequency sweep profile of nuclear enhancement with radiation dose we hypothesise may be due to some degree of cooperativity (interactions) between electrons when the radical concentration becomes sufficiently high that the picture of an isolated, well-defined radical species as an independent molecule breaks down. The temperature-dependent electronic $g$-shift we observed in the EPR data is consistent with magnetic ordering in our system, or at the very least some degree of electronic reconfiguration over molecular length scales (c.f. Fig. S6), leading to a superposition of effective radical species that may interconvert and interact with neighbouring molecules (alanine has four in the unit cell), including the primary electronic radical disfavoured at high temperature. Indeed, spontaneous spatial ordering in neutral alanine crystallites has previously been reported to occur in strong magnetic fields, (?) arising out of magnetic anisotropy of the L-alanine unit cell and its comparatively large (in magnitude) $\chi_{m}$; (?, ?, ?) and zirconium-nitrate doped $L$-alanine single crystals form magnetic order and have been used as nonlinear optical devices. (?) While traditional paramagnetic systems maintain independent spin behaviour and molecular character even at low temperatures, our partially ordered alanine crystallites dispersed in glycerol could exhibit more complex magnetic phenomena as radical centres become close enough to develop short-range magnetic correlations at low temperatures and high radiation doses within individual crystallites. The glycerol matrix in this scenario could function analogously to domain boundaries in ferromagnetic materials. By physically separating the alanine crystallites, glycerol creates discontinuities in the exchange pathways, limiting the correlation length of any magnetic ordering, but permitting thermodynamically driven spin diffusion through weak coupling provided by proton/carbon spins in the matrix. This is conceptually similar to the role of Bloch walls in ferromagnetic materials, though operating at the nano- or microscale between distinct crystallite domains rather than within a continuous bulk material. This model could potentially explain several of our observations:
Firstly, the spatial confinement of magnetic correlations to individual crystallites would naturally lead to anisotropic magnetic responses dependent on crystallite orientation, consistent with our texture analysis showing preferential alignment. Each crystallite may develop its own local magnetic environment influenced by its specific orientation relative to the external field.
Secondly, this partial magnetic ordering could modify the effective field experienced by nuclear spins, altering the resonance conditions for polarisation transfer for that microscopic environment. In conventional DNP models, polarisation transfer is mediated by electron-nuclear dipolar interactions under well-defined frequency matching conditions (e.g. $\omega_{e}\pm\omega_{N}$ for the solid effect) and the angular orientation terms in the dipolar interaction Hamiltonian ($3\cos^{2}\theta-1$) are assumed to be spatially averaged in a glass. This regionality in orientation together with a transition from individual to collective behaviour would explain why the best conventional theoretical models of DNP fail to predict our observations at higher radical concentrations, as they assume only electron-nuclear interactions and/or are based on thermodynamic arguments assuming some degree of spatial homogeneity and good thermal contact. They do not account for many-body effects that emerge when multiple electrons (arising from multiple radicals in close proximity) interact collectively.
Thirdly, the glycerol “boundaries” could create a heterogeneous distribution of magnetic environments throughout the sample. Radicals near the glycerol interface might experience different effective fields than those in crystallite interiors, leading to a distribution of resonance conditions that broadens and shifts the optimal DNP frequency. This domain-like structure might also explain the effectiveness of frequency-swept microwave pulses in our system. By sweeping across a range of frequencies, we could be sequentially addressing different sub-populations of radicals with distinct magnetic environments, each contributing to the overall polarisation enhancement. The constructive summation of these contributions could explain the superior performance compared to fixed-frequency irradiation, even though the frequency swept range is small compared to the overall EPR lineshape. It also provides a mechanism for the biexponential growth of polarisation, which could be ascribed as being that within each crystallite occurring quickly, and that between them occurring through the dilute matrix of natural abundance carbon-13 spins in glycerol. These results are overall surprising, and future work will help elucidate a mechanism behind them. They represent a new paradigm in DNP, where a partial spatial ordering of spin systems creates emergent behaviour not predicted by isolated electron models of molecular physics.
The levels of nuclear enhancement we achieved ($P\approx 17\%$ or $\epsilon\approx 142\,000$ at 1.4T) compare favourably to the $\epsilon\approx 700$ to 800-fold enhancements reported at 11.7T for ionising-radiation radical induced DNP previously by Co-60 $\gamma$-rays by Giannoulis, (?) and are comparable with the optimal trityl, narrow line-width radicals used clinically. For a non-irradiated alanine sample prepared with trityl radical, the polarisation level at the point of dissolution was 19.2% and the nuclear T1 was 40.8 s at 1.4T. In contrast to the work by Giannoulis, we did not observe data consistent with a simple model by the solid-effect, and obtained qualitatively different frequency-sweep curves. This may be due to differences in the radiation transport physics of ultra-high dose rate 6 MeV electrons vs $\sim$1 MeV $\gamma$-rays, which will deposit more charge into the crystallites, as well as dose rate effects altering the environment around the radical centre, by, for example, introducing a degree of unresolved crystallographic strain. (?, ?, ?) We also note that Giannoulis’ alanine samples were flame-sealed and irradiated in vacuo (our samples were in air), and also polarised as dry powders. Finally, for reference, a previous study where alanine was also hyperpolarised with the trityl radical yielded a liquid state polarisation of 12.6% at the point of dissolution and an in vitro T1 of 41.5 s at 3 T (?). Our irradiated sample produced comparable results; polarisation levels of 17.7% at the point of dissolution were achieved, while the extracted nuclear T1 relaxation times were 77.3 s at 1.4 T. The nuclear T1 relaxation time was prolonged due to the fact that the irradiation-generated radicals quench upon dissolution. Future work could explore the direct deuteration of the 13C-enriched alanine to potentially enhance polarisation levels as previously observed for other molecules hyperpolarised by dDNP (?). Crucially, the pH of the irradiated alanine samples (approximately 7) also made them suitable for in vivo trial with dissolution in water; we note that neutral salts are readily crystallised and substances such as sodium pyruvate likely will be readily amenable to this technique.
Beyond the physical novelty of our approach, our in vivo experiments demonstrate significant translational potential. Electron-irradiated 13C-alanine not only matched but in some aspects exceeded the performance of conventional preparations using exogenous radicals. It is worth considering how 13C-alanine could compare to 13C-pyruvate, the most commonly used metabolic probe with dDNP and undergoing clinical trials. Whilst pyruvate polarises very well with dDNP and is rapidly metabolised to lactate, alanine, and bicarbonate, it has been noted that the pyruvate/lactate ratio is not reflective of intracellular redox state following HP pyruvate infusion (?). Upon switching the injectable substrate from 13C-pyruvate to 13C-alanine, Hu et al. recorded more than a ten-fold increase in the 13C-lactate/13C-pyruvate ratio (?). The difference was suggested to be due to the uptake of 13C-pyruvate being limited by perfusion and downstream flux through monocarboxylate transporters during the acquisition window. An incomplete uptake of 13C-pyruvate would mean contributions from an extracellular 13C-pyruvate pool in the measured 13C-pyruvate signal. Furthermore, any extracellular 13C-pyruvate present in plasma can be converted into 13C-lactate by red blood cells that contain lactate dehydrogenase, leading to an “extracellular” 13C-lactate pool (?). Signals arising from intracellular and extracellular pools cannot be easily distinguished in hyperpolarised MR, leading to the proposed use of 13C-alanine for determining the 13C-lactate/13C-pyruvate ratio. For 13C-alanine to be converted into 13C-pyruvate, and subsequently 13C-lactate, it must first be exposed to the intracellular enzyme alanine transaminase. Thus, the 13C-pyruvate signal detected following the injection of 13C-alanine would be intracellular in origin. Moreover, after the metabolism of alanine is observed and quantified, the cellular redox potential can be easily determined  (?) from the known relation $K_{\text{eq}}=\frac{[\text{Pyruvate}]\,[\text{NADH}][\text{H}^{+}]}{[\text{Lactate}][\text{NAD}^{+}]}$ where $K_{\text{eq}}=1.11\times 10^{-11}$ M. If a measure or estimate of intracellular pH is obtained – for example, from 31P-MRS (?) or through the use of hyperpolarised pyruvate (?) – this would provide a direct mechanism by which the NAD/NADH ratio could be non-invasively determined, as has been reported previously for trityl-hyperpolarised alanine. (?)
Our electron irradiation approach uniquely combines this biochemical advantage with the practical benefits of in situ radical generation and spontaneous quenching, providing a powerful new tool for quantifying intracellular redox potential. It is also the first to explicitly consider partially ordered solids as substrates for biomedical dDNP rather than amorphous glasses that are commonly used with the technique (and, indeed, often listed as a requirement for its success;  (?)) While this technique could be extended to other substrates, 13C-alanine represents an ideal initial candidate due to its metabolic significance, detailed quantum-mechanical study, and demonstrated biochemical and clinical relevance. This approach may reduce the technological, regulatory, and economic barriers that have constrained hyperpolarised MRI primarily to specialised research centres, potentially expanding access to this powerful diagnostic capability.
3.1 Limitations
We should note that several limitations should be acknowledged in this work: notably, our mechanistic understanding of the DNP process in this system remains incomplete and in need of extension via theoretical modelling. The unusual frequency sweep profiles of this highly efficient polarisation process hint at cooperativity and physics beyond conventional DNP models, but a comprehensive theoretical framework that quantitatively predicts these behaviors is lacking. Additional spectroscopic studies with variable temperature, field strength, and radical concentration would help elucidate the underlying mechanisms.
Additionally, while we demonstrated stable radical formation in alanine, the applicability of this approach to other metabolically relevant molecules requires further investigation. Different molecular structures may respond differently to electron irradiation, potentially generating distinct radical species with varying stability and DNP efficiency. The specific arrangement of alanine in a crystalline lattice likely contributes to the observed radical stability and its partial ordering in glycerol (analogous to a nemantic liquid crystal), and molecules with different crystal structures, or those that do not readily crystallise, may present challenges. Furthermore, the comparatively high radiation doses considered may embrittle plastics used in the construction of fluid paths, but this phenomenon is well-characterised, and many medical plastics are routinely sterilised with comparable doses. (?, ?) We chose to investigate alanine because it is exceptionally well studied from a radiation physics and EPR point of view, and is a reported biomolecular probe investigated for probing central metabolism. Other, perhaps more commonly employed biomolecules may well be amenable to this technique, where significant further study across all aspects of this highly interdisciplinary science remain yet to be elucidated. We can propose three fundamentally distinct approaches to explore this in future work: (1) by directly irradiating other molecules; (2) by irradiating alanine powder containing microcrystallites (such as that here) and dispersing alanine radicals in the solid state with other molecules of interest; or (3) by utilising co-crystallised (e.g. natural abundance) alanine and an isotopically labelled molecule of interest, under the assumption that radical species formed act as dilute spin impurities in the resulting system. We have yet to explore any of these other exciting avenues for the technique.
Finally, the highest nuclear polarisation levels achieved with our approach ($\sim 20\%$) remain lower than those reported for optimised pyruvic acid preparations with trityl radicals ($50\%$). While sufficient for in vivo imaging, higher polarisation would improve SNR and potentially allow detection of lower-concentration metabolites. Further optimisation of irradiation parameters, alanine crystallite size, and DNP conditions might narrow this gap. Likewise, owing in part to this limitation, our in vivo experiments were limited to proof-of-concept demonstrations in two healthy animals, precluding a detailed characterisation of the technical vs biological variability of the technique beyond providing proof-of-principle. The performance of irradiated alanine in disease models, particularly its sensitivity to pathological alterations in redox potential, requires validation in relevant preclinical models before clinical translation. We also note that the applicability of this technique may well be higher in other diseases and organs where alanine metabolism is believed to be more avid; for example, in the liver where the lactate/pyruvate ratio is closer to 10 following the infusion of alanine (rising to 20 after an ethanol challenge). (?)
Beyond scientific limitations, there is much that remains to be proven for the clinical adoption of this technique. At present, most European sites conducting human hyperpolarised experiments use electron-beam irradiation to sterilise (25-45 kGy) empty ‘sterile fluid paths’ made of polyethyletherketone (PEEK) and polyphenylsulfone (PPSU) plastics, which are filled under sterile conditions by a pharmacy team with pyruvic acid/trityl radical, snap frozen in liquid nitrogen, purged with helium under pressure, laser welded, and transported to the polariser for ultimate human use. (?) We envisage that filling a fluid path with reagents, irradiating the solid isotopically labelled compound to a higher level and the fluid path overall to a level for sterilisation would be a feasible method to generate transportable, sealed single-use devices in a centralised “worldwide” factory with an electron accelerator. However, we have not explicitly demonstrated this, nor have we shown that PEEK or PPSU walled tubes could withstand the high radiation doses required. We likewise have not undertaken the detailed preclinical toxicity analyses to show that no byproducts or adjuncts would be formed by high energy electron irradiation of PEEK or PPSU plastics containing solid isotopically labelled crystallite alanine, all of which is future work and necessitate completion in a validated, accredited laboratory prior to human use.
4 Materials and Methods
4.1 Sample Preparation
The samples used in this study can be categorised into three main groups: alanine containing endogenous irradiation-generated radicals, alanine mixed with the exogenous trityl radical OX063 at a concentration of 38 mM, and pyruvate with 30 mM AH1111501. These exogeneous radicals differ only by an exterior methyl ‘R’-group in their chemical structure and have almost identical electronic properties; these concentrations of both trityl radicals were chosen to reflect reported optimal conditions for DNP for both molecules.
[…]
4.2 Dynamic Nuclear Polarisation
Samples were polarised at 3.35 T or 6.7 T using a HyperSense (Oxford Instruments) or SpinAligner (Polarize), respectively. Further details are provided in the SI §6.2.
4.3 In Vitro Experiments
The polarisation levels and nuclear T1 relaxation times of samples were recorded using a benchtop NMR spectrometer at 1.4 T (SpinSolve 60ULTRA Carbon, Magritek). The accompanying software also enabled the level of polarisation to be estimated at the point of dissolution.
4.4 In Vivo Experiments
Two healthy male Sprague-Dawley rats were injected with one shot of each type of hyperpolarised alanine sample via a tail vein catheter. The time between each injection was approximately two hours. The respiratory rate and body temperature of each rat were monitored using an MRI compatible small-animal monitoring system (Small Animal Instruments Inc, USA) whilst the rats were under anaesthesia with 2.5-3% sevoflurane in 1 l/min medical air. Both rats were scanned using a dual-tuned 13C/1H volume rat coil at 3 T (GE MR750, GE Healthcare). One rat was scanned using a slice-selective spectroscopic acquisition, the other imaged using a 2D MR Spectroscopic Imaging (MRSI) sequence, EPSI, using fidall. (?)
Prior to these scans, a urea phantom was used to calibrate the applied flip angles, and proton images were acquired for anatomical reference. For the slice-selective spectroscopic acquisition, 80 mm slice thickness was used to ensure the whole abdomen of the rat was imaged. 1 ml of hyperpolarised sample was injected over approximately 5 s, and the scan started upon the start of the injection. A 30  flip angle was applied with a repetition time (TR) of 3 s for a total of 5 minutes (100 time points). For the MRSI, 1 ml of hyperpolarised substrate was injected over approximately 5 s, and the scan started upon the start of the injection. A $16\times 16$ matrix was chosen, and a 15  flip angle was applied with a TR of 64 ms for a total of 2 minutes (8 time points and approximately 16 s per time point). The pH of the injected solutions was measured to ensure physiological compatibility, with values ranging from 6.7 to 7.4, well within the acceptable range for intravenous administration. All animal experiments were conducted following appropriate independent ethical review and under appropriate licensing regimes in Denmark.
4.5 Electron Paramagnetic Resonance
An X-band continuous wave (CW) spectrometer (EMXmicro, Bruker BioSpin; Super High Sensitivity rectangular cavity probe head) was used to record EPR spectra at room temperature. For recording spectra at 5 K and measurement of $T_{1e}$ and $T_{2}$, an X-band pulsed spectrometer (E680, Bruker BioSpin) was used. The Chromium(III) impurity in Magnesium Oxide powder, a common standard in EPR, (?, ?) was used as a frequency reference. Further details are provided in SI §6.7.
Samples that comprised dry powder were directly loaded into 4 mm thin wall quartz EPR tubes, and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25–35 . Liquid state samples were first loaded in 2 mm diameter capillary tubes, which in turn were loaded into the EPR tubes.
Spin counts were determined from EPR spectra using an in-house developed MATLAB script. The script incorporated baseline correction, normalised for differences in resonator $Q$-factor, and integrated the recorded spectra.
13C spectra were fitted via a constrained Bayesian optimisation routine (?). The line widths and shapes of spectra were also compared to models built with the ‘pepper’ function in EasySpin, a validated QM spectral simulation library designed for EPR. (?)
4.6 X-ray Diffraction
XRD was conducted using copper $K_{\alpha}$ X-rays, of wavelength 1.5405982 Å and 1.544497 Å, in an Empyrean (Malvern Panalytical) diffractometer. XRD samples comprised dry alanine powder or an alanine powder/glycerol mix (1:1 w/w). The Cu $K_{\beta}$ wavelength was removed from the diffraction signal with a Ni filter placed in front of the PIXcel1D detector. A set of programmable divergence slits were used for the incident optic, set to illuminate a fixed surface area of 1 cm2 across the angular range of the scan. During the scans, the samples were rotated about the azimuthal axis at a frequency of 0.5 Hz to fully characterise the crystallite distribution.
XRD data were compared to a previously reported crystal structure of alanine, obtained via neutron scattering (?), and a Rietveld refinement was undertaken in Maud using a spherical harmonic texture basis (?). Further details are provided in the SI §6.9. [7890 Words]
5 Figures and Tables
Figure 1: Radical generation and quantification. (A) An ultra-high dose rate optimised in-house developed linear accelerator was used to bombard (B) a target containing polycrystalline [1-13C]alanine with 6 MeV electrons. (C) These generate Stable Alanine Radicals (SAR) detectable by EPR in a dose-dependent fashion compared to (D) TEMPO as a concentration standard. (E) These EPR data were acquired in a single CW EPR scan at 9.38611 GHz with a modulation frequency of 100 kHz; and sweep rate of 0.816 mT/s. The 13C-hyperfine interaction in the SARs, significantly alters the EPR spectrum compared to natural abundance (NA) alanine, which we modelled (F) via simulations in EasySpin, shown frequency-corrected to a standard frequency of 9.6 GHz, with our 13C-model in black (detailed in SI §6.8).
Figure 2: DNP and Glycerol addition. (A) DNP build-up profiles of 13C-NMR signal arising from different high dose rate irradiated samples at 3.35 T; we found that dispersing irradiated polycrystalline alanine powder in glycerol produced substantially higher polarisation. (B) Longitudinal measurements of the spin counts showed that there was no significant effect on the endogenous radical concentration with increased duration of storage in anhydrous glycerol up to 4 months, in sharp contrast to pyruvic acid/trityl radical mixtures which quench within hours (mean $\pm$ SD of three technical replicates). (C) Alanine/glycerol mixtures did not significantly alter acquired EPR data at room temperature (RT), but at (D) 5 K the electronic environment altered considerably, and the EPR spectrum recorded shows significant changes from RT data. These data were obtained via either CW EPR as previously (RT) at 9.38611 GHz or at 5 K on a pulsed instrument at 9.76344 GHz and have been frequency-corrected to 9.6 GHz in (D) or 9.4 GHz in (C), for display on a common scale; c.f. SI §6.7.
Figure 3: DNP at 6.7 T and partial order(A) The frequency sweep profile of DNP with alanine/glycerol mixtures showed qualitative differences with increasing radiation dose, which we hypothesise is due to cooperativity. (B) Polarisation build-up as a function of time showed profound biexponential behaviour under these conditions, distinct from narrow linewidth radicals such as trityl; all arising from 13C-NMR on [1-13C]alanine. (C, Di) X-ray diffractometery measurements of both samples were able to demonstrate that polycrystalline alanine remained dispersed (not dissolved) in glycerol and was well described by (Dii) existing crystal structures with slight refinement but had profound crystallographic texture, indicative of an ordered crystalline arrangement within the glycerol matrix. (E) The efficiency of DNP under these conditions is surprising, comparable to narrow linewidth trityl radicals. Further experimental details are provided in the SI.
Figure 4: Polarisation optimisation and high-field comparisons with trityl. (A) At 6.7 T, a unimodal, always-positive frequency sweep curve was observed. Pulsed modulation with a swept-frequency microwave pulse as illustrated obtained a significant increase in nuclear polarisation. (B) The molecular orientation of the alanine radical matters significantly for the spectral density function $g(\omega)$ as defined by Wenckebach, (?) simulated here using our model derived from powder 13C alanine EPR data and assuming a partial ordering parameter $\lambda$. (C) The resultant limiting polarisation obtained was high, and comparable to the use of trityl radicals with alanine, with biexponential fit shown. (D) After dissolution, spontaneous quenching of endogeneous radicals formed by electron irradiation resulted in a lengthened $T_{1}$ (77 s) compared to that with trityl (40 s) at 1.4 T (fits shown in C). Further experimental details are provided in the SI.
Figure 5: In vivo demonstration in the rat. (A, B) Stacked and (C, D) summed spectra from slice-selective spectroscopic acquisitions following separate tail-vein injections of [1-13C]alanine polarised via 38 mM OX063 trityl radical and hyperpolarised irradiated 13C-alanine. (C, D) Inset: a magnification of the spectra presented in the panels to highlight the intensities and positions of the downstream metabolites, whose chemical shift values match those reported in the literature for 13C-lactate and 13C-pyruvate. (E, F) After quantification by AMARES, metabolite dynamics were comparable between the two methods of preparation. (G) MR Spectroscopic Imaging (MRSI) images obtained in a separate rat were likewise near identical, validating the use of the technique for in vivo imaging of redox potential, alanine images of which are shown concatenated over time. Further experimental details are provided in the SI.
Table 1: Summary of DNP build-up parameters from biexponential fits across different experimental conditions and radical sources. Build-up profiles follow the biexponential model: $y=P[A(1-e^{-t/T})+(1-A)(1-e^{-t/\tau})]$, where $P$ is the build-up constant, AA is the fast component weighting, $T$ is the fast component time constant, and $\tau$ is the slow component time constant. It was not possible to fit 3.35 T data to a biexponential curve stably; and likewise a biexponential fit to trityl data at 6.7 T returned monoexponential behaviour. Arbitrary units are comparable within experiments but not between them.
Fig. Sample/Condition Dose Build-up Fast Slow
(kGy) Amplitude Time $T$ Time $\tau$ Weighting AA
$P$ (a.u.) (s) (h)
2a 10 kGy (glycerol plug) 10 $9.59\times 10^{2}\dagger$ Monoexponential fit poor; $\tau_{\text{mono}}=3171342310$ s
100 kGy (glycerol plug) 100 $6.28\times 10^{3}\dagger$ ($\tau_{\text{mono}}=19236$ s)
100 kGy (glycerol mix) 100 $2.37\times 10^{4}\dagger$ ($\tau_{\text{mono}}=2836$ s)
3b 30 kGy (glycerol mix) 30 $1.52\times 10^{6}$ $22.9$ $0.92$ $0.264$
50 kGy (glycerol mix) 50 $8\times 10^{6} 22.22$2.2 $2.69$ $0.193$
70 kGy (glycerol mix) 70 $5.11\times 10^{6}$ $14.9$ $6.88$ $0.099$
[…]
70 kGy (glycerol mix) 70 $1.7\times 10^{1}$ 2020 $0.481$ $0.088$
Notes: $\dagger$ These data are at 3.35 T.
$\ddagger$ OX063 trityl radical shows purely monoexponential behaviour ($A=0$ in one fit; identical time constants in the other).
Acknowledgments
Funding:
We would like to acknowledge support from the Novo Nordisk Foundation (NNF21OC0068683); the British Heart Foundation (refs. RG/11/9/28921, RE/13/1/30181, FS/19/18/34252, and FS/14/17/30634); the Wellcome Trust (221805/Z/20/Z); the Engineering and Physical Sciences Research Council (EPSRC) and Medical Research Council (MRC) (EP/L016052/1); and the Colleges of St. Hugh’s and Somerville in the University of Oxford. We would furthermore like to acknowledge the University of Oxford British Heart Foundation Centre for Research Excellence (RE/13/1/30181) and the NHS National Institute for Health Research Oxford Biomedical Research Centre programme. The views expressed are those of the authors and not necessarily those of the NIHR or the Department of Health and Social Care. AMB is grateful to The Royal Society and the EPSRC for a Dorothy Hodgkin Fellowship (DH160004 and DHF/R/221018), and the University of Manchester for a Dame Kathleen Ollerenshaw Fellowship. AMB also thanks the Royal Society of Chemistry for a Community for Analytical and Measurement Science fellowship (CAMS Fellowship 2020 ACTF ref 600310/09). AMB is grateful for support from National Research Facility for EPR (EP/W014521/1, NS/A000055/1, EP/V035231/1, and EP/S033181/1). We also gratefully acknowledge Cambridge Isotope Laboratories, Inc, who provided a gift of [1-13C]alanine to JJM and Z.R. We also thank Christiane Timmel for useful discussions about quantitative EPR spectroscopy. JJM would additionally like to thank Professor Elspeth Garman and Professor Ian Carmichael for both dinner and many useful discussions about radiation chemistry in crystals. The practical assistance of the University of Oxford Department of Physics Mechanical Workshops is also kindly noted.
Author contributions:
CHER and JYCL conceptualised research, conducted experiments, analysed data, and drafted the manuscript. ESSH, NVC, and DAD assisted with sample preparation and in vivo/high field DNP experiments; BWC and SS assisted with low-field DNP experiments. KP, IT, BV, CHER undertook electron linear accelerator experiments for sample irradiation, including design and manufacture of equipment. JL, CHER and JJM undertook X-ray diffraction experiments and their interpretation. WM, and AMB provided expertise in EPR spectroscopy and data interpretation. ZR assisted with modelling. LBB and CL supervised the in vivo experiments and contributed to experimental design. JYCL, DJT and JJM conceived and supervised the project, secured funding, analysed data, and finalised the manuscript. All authors reviewed, edited and approved the final manuscript.
Competing interests:
JYCL, CL, JJM, CHER and DT are listed as authors on a filed patent pending examination (EP24168322.6) pertaining to the use of ultra-high dose rate electron irradiation for preparing hyperpolarised contrast agents. All other authors have no conflicts of interest to declare.
Data and materials availability:
Crystallographic data pertaining to this project are enclosed as CIF files as SI to this paper, as is EPR raw data. Materials (i.e. ultra-high dose rate irradiated labelled metabolites) are available to others upon the signing of a collaboration and material transfer agreement. All other data are available in the main text or the supplementary materials.
Supplementary materials
Materials and Methods
Supplementary Text
Figs. S1 to S9
Tables S1 to S2
References (123-0)
Data S1 to S2
Supplementary Materials for
Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation
Catriona H. E. Rooney, Justin Y. C. Lau, Esben S. S. Hansen, Nichlas Vous Christensen, Duy A. Dang, Kristoffer Petersson, Iain D. C. Tullis, Borivoj Vojnovic, Sean Smart, William Myers, Zoe Richardson, Jarrod Lewis, Brett W. C. Kennedy, Alice M. Bowen, Lotte Bonde Bertelsen, Christoffer Laustsen, Damian J. Tyler, and Jack J. Miller∗‡
Corresponding author. Email: jack.miller@physics.org $^{\dagger},\ddaggerThese$ authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S9
Table S1
Other Supplementary Materials for this manuscript:
Data S1 (Crystallographic)
Data S2 (EPR)
6 Supplemental Experimental Methods
6.1 Irradiation
Full technical details of the linac are published elsewhere and much of its construction details are freely available at https://users.ox.ac.uk/~atdgroup/technicalnotes/.
Renderings of the sample holder used to hold the sample during the irradiation are shown Fig. S1A. To produce the samples, a block of aluminium with two sample spaces was loaded with dry alanine powder. The powder was secured in place by taping two ordinary glass microscope slides either side of the aluminium block. The glass-aluminium-glass sandwich was held in an upright position by microscope slide spring clips within another aluminium component that itself was connected to a high precision linear slider. The integration of the slider and two stoppers into the design facilitated the centring of each sample space within the beam path. To increase the homogeneity of the irradiation exposure across the total volume of each sample space, the glass-aluminium-glass sandwich was flipped 180° after half of the total desired irradiation dose had been delivered. Irradiation doses over the range 10–100 kGy were applied to produce samples with a range of different endogenous radical concentrations.
A radiation transport simulation was undertaken in TOPAS, (?, ?) an open-source radiation transport code optimised for medical physics applications based on Geant4, written by CERN. (?) As shown in Fig. S1B, C, this was used to optimise dose distribution across the sample with regards to the ‘pile up’ dose deposited either in the glass coverslips or aluminium target; the preferred deposition function is that of a ‘top hat’ within the sample but it was found that appropriate spacing and modulation of design produced a reasonable distribution function (Fig. S1D). A significant amount of charge transfer is expected to take place within the electron bombardment process. In order to obtain a more homogeneous dose distribution within the sample, we rotated the sample by 180  in the middle of the irradiation process; after irradiation, the sample was mixed by hand, further homogenising it.
Figure S1: Details of the overall process of e-beam irradiation. A: A sample holder for the irradiation of alanine samples using a 6 MeV high-dose rate electron linear accelerator was designed. Renderings of the sample holder comprising (1) an aluminium endplate with a 5 cm diameter aperture for the electron beam, (2) a rectangular block of aluminium — with two cylindrical holes each acting as a sample space — interposed between two glass microscope slides and secured upright with microscope slide spring clips to (3) an L-shaped aluminium block attached onto (4) a high precision linear slider. (5) Two stoppers were positioned either side of the slider to facilitate the positioning of each sample. B, C: Simulated Monte-Carlo radiation transport characteristics of the designed target demonstrating appropriate dose deposition inside the compressed alanine pellet (red denotes leptons; green photons); together with D: predicted dose deposition as a function of depth, normalised such that the total dose is unity.
6.2 DNP
DNP was performed either using an Oxford Instruments HyperSense polariser at 3.35 T or a customised SpinAligner 6.7 T polariser. In both cases either microwave sweep profiles were obtained by iterating over microwave irradiation frequency, or polarisation build-up curves in which a low flip angle pulse was repeatedly played and polarisation monitored. The carbon Larmor frequency at both field strengths were 35.96048861 MHz and 71.816528 MHz respectively. Amplitudes obtained and reported as signal during microwave sweeps or build-up curves are integrated over the well-defined NMR peak present in the 13C NMR signal as is commonly performed with DNP.
Reported time-constants for the DNP build-up profiles were fitted to a biexponential curve given by the equation
$y=P\left(A\left(1-e^{-\frac{\text{time}}{T}}\right)+\left(1-A\right)\left(1-e^{-\frac{\text{time}}{\tau}}\right)\right)$ (S1)
where $P$ is the build-up constant, $A$ and 1A1-A are the relative weightings of the fast and slow components of the polarisation build-up, and $T$ and $\tau$ are the fast and slow components of the build-up times.
For the HyperSense, microwave frequency sweeps were acquired for each sample between 93.75 to 94.195 GHz in 10 MHz steps with a two minute build-up per frequency step. The microwave power was held constant at 100 mW for all microwave frequency sweeps and build-ups. The build-up of polarisation was subsequently monitored at the peak frequency of the microwave frequency sweep every three minutes using a low flip-angle readout. A 3.35 T frequency sweep is shown in Figure S2 together with the corresponding sweep for an alanine + OX063 sample. A small potential enhancement lobe may be seen at 93.970 GHz. Black dashed lines denote equivalent features in the 6.7 T microwave sweep curves, scaled exactly for field. Given the considerably less sophisticated microwave and RF architecture of the HyperSense compared to the SpinAligner, the possibility that this lobe at 93.970 GHz is artefactual remains, although it was not observed in the control sample.
Figure S2: The frequency sweep obtained at 3.35 T for both alanine and trityl radical and 100 kGy irradiated alanine. The parameters noted “LL” (lower limit), “UL” (upper limit) and “COM” (centre of mass) are presented for illustrative comparisons with higher field DNP data, and have been defined phenomenologically as the lower, centre and upper region of features in the 6.7 T curve scaled for the exact difference, i.e. the corresponding down-scaled values of [187.77, 187.95, 188.15][187.77,\,187.95,\,188.15] GHz at 3.358 T.
Based on previous optimisation work and the use of a narrow line width radical, the SpinAligner was used without microwave modulation for the non-irradiated samples. For the irradiated samples, microwave modulation was applied over a range of 25 MHz about the given centre frequency, sinusoidally varied with the frequency of microwave irradiation occurring at a rate of 1 kHz. This applies a weak modulation and attempts to invert a greater proportion of electron spins; it also is recommended by the microwave source manufacturer in order to promote greater thermal stability of their system.
[…]
The NMR spectra recorded during microwave frequency sweeps and polarisation build-ups were quantified by peak integration in MATLAB (ver. 2023a; The Mathworks, Inc). Polarisation estimates were obtained in the liquid state both hyperpolarised and at thermal equilibrium, and then back-calculated using the measured $T_{1}$ and known time of dissolution to estimate solid-state polarisation; for samples with low nuclear polarisation in the solid state (where relative error post dissolution is likely higher or sufficiently high to be inaccurate) a direct linear scaling of ADC values obtained at constant gain with a known polarisation reference were used.
6.3 Power characterisation at 6.7 T
Figure S3: (A) Frequency sweep profiles comparing 10 mW and 20 mW microwave power showing preserved asymmetric enhancement, and marginally enhanced build-up. (B) Build-up profiles demonstrating increased absolute signal at higher power with largely unchanged kinetics. (C–G) Systematic investigation of frequency modulation parameters showing optimal conditions near 100 Hz modulation rate with preserved biexponential build-up characteristics across all conditions tested.
To address potential concerns regarding the optimisation of microwave irradiation conditions, we performed extensive calibration experiments investigating the effects of microwave power, and, for our source in the SpinAligner, frequency modulation and pulse timing parameters on the DNP enhancement profiles. These experiments were essential given the high-field (6.7 T, 188 GHz) operation of our system, where microwave field inhomogeneity within the sample cavity becomes increasingly problematic (?). This is because the SpinAligner’s microwave cavity (diameter 28 mm, height 30 mm) is highly overmoded at 188 GHz, precluding meaningful quality factor measurements that may come with unequal power deposition into the sample.
Additionally, owing both to the closed-cycle cryogenic nature of the machine and additional hardware constraints, continuous-wave irradiation was not feasible due to thermal instabilities in both the microwave source and the sample environment. This necessitated the pulsed modulation approach, which while effective for our current studies, should not be confused with (and, indeed, precludes) the investigation of coherent polarisation transfer mechanisms such as TOP-DNP (?) or NOVEL sequences (?, ?) that might further optimise the enhancement – our modulations are orders of magnitude slower than the electron Larmor frequency.
Systematic power-dependent measurements were performed on 70 kGy electron-irradiated 13C-alanine samples using our “standard conditions” that were used in the majority of this work (25 MHz amplitude modulation, 100 Hz frequency modulation, both detailed subsequently). Figure S3A shows frequency sweep profiles acquired at 10 mW and 20 mW output power, demonstrating that the characteristic asymmetric enhancement profile is preserved across this power range. The build-up kinetics (Figure S3B) show increased absolute signal intensity at higher power, consistent with more efficient microwave saturation of the electron spin transitions, but the fundamental biexponential character and relative time constants remain unchanged.
The SpinAligner system employs pulsed microwave irradiation with frequency modulation to accommodate the broad EPR linewidth of the stable alanine radicals while avoiding continuous-wave heating limitations of the 188 GHz source. The user-specified parameters are detailed below:
  • AM (Amplitude Modulation): The frequency sweep bandwidth in MHz. This determines the range of microwave frequencies scanned during each pulse (e.g., 25 MHz spans $\pm$12.5 MHz around the center frequency).
  • FM (Frequency Modulation): The base modulation rate in Hz. This parameter, combined with hardware duty cycle constraints, determines the pulse repetition timing.
The relationship between these user parameters and actual pulse timing is governed by the microwave synthesiser hardware implementation. For combined PM+FM operation, the pulse timing follows:
$t_{\text{pulse on}}$ $=t_{\text{off,displayed}}+\frac{10^{6}~\mu s/s}{f_{\text{FM}}}-40~\mu s$ (S2)
$t_{\text{pulse off}}$ $=t_{\text{on,displayed}}+40~\mu s$ (S3)
where $f_{\text{FM}}$ is the FM parameter in Hz, and the displayed on/off times are hardware-inverted values read from the synthesiser. The term $10^{6}/f_{\text{FM}}$ represents the base period in microseconds corresponding to the FM rate.
For our standard conditions (AM = 25 MHz, FM = 100 Hz), these equations yield:
$t_{\text{pulse on}}$ $=\frac{10^{6}}{100}-40=9960~\mu s\approx 9963~\mu s\text{ (measured)}$ (S4)
$t_{\text{pulse off}}$ $=48139~\mu s\text{ (measured)}$ (S5)
This corresponds to a duty cycle of approximately 17% and a total cycle time of  58 ms. The discrepancy between the theoretical base period (10 ms for 100 Hz) and actual cycle time reflects hardware-imposed timing constraints and thermal management requirements.
During each “on” period, the microwave frequency is swept linearly across the AM bandwidth (e.g. 25 MHz) in discrete steps. The number of frequency steps and dwell time per step are determined by the synthesiser’s internal timing, typically resulting in 25–50 discrete frequency points per sweep. This approach aims to ensure that the broad EPR linewidth ($\sim$15 mT FWHM, corresponding to $\sim$420 MHz at 6.7 T) is broadly saturated while maintaining sufficient power at each frequency point to drive polarisation transfer.
We systematically investigated the effects of both amplitude modulation bandwidth (25–100 MHz) and frequency modulation rate (100 Hz to 10 kHz) on the enhancement characteristics. Figure S3C–G summarises the parametric dependence of the biexponential build-up characteristics on modulation frequency. Although very preliminary values and potentially specific to our hardware, the build-up constant $P$ (panel C) shows optimal values near 100 Hz modulation, with diminishing returns at higher frequencies. This behavior likely reflects the interplay between the EPR linewidth ($\sim$15 mT FWHM), the frequency sweep range (25 MHz), and the electron spin relaxation times under our experimental conditions.
Importantly, the relative weightings of fast and slow build-up components (panels D and E) and their respective time constants (panels F and G) show some systematic but modest variations with modulation parameters. The fast component time constant $T$ remains in the range 300–400 s across all conditions tested, while the slow component $t$ varies between 0.5–2 hours (the upper end of the build-up experiment tested). We believe this rules out, for example, a finite microwave bandwidth leading to differential enhancement, and we further note that the pyruvic acid frequency sweep curve under these conditions is a well-resolved bimodal curve of equal intensity within this region.
6.4 Stability analysis
Samples of irradiated alanine and glycerol were created on or the day after irradiation, and stored in borosilicate glass haematocrit tubes, themselves stored in a falcon tube. These were stored in separate home-made desiccators consisting of monolayer of silica desiccant gel beads (Sigma Aldrich, UK) in a plastic ‘tupperware’ self-sealing container. Three repeated EPR measurements were obtained by inserting the haematocrit tubes inside an EPR tube, and were rotated by hand within the machine between acquisitions. EPR spectra were obtained at weeks 0, 2, 4, and 16 after irradiation. A simple linear regression approach was used for 30, 50 and 70 kGy irradiated samples dispersed in alanine; there was no significant change with time ($p$ = 0.14, 0.085, 0.18, respectively) and the goodness of fit was poor ($R^{2}=0.20$, $0.27$, and $0.17$ respectively). We interpret this as providing evidence of stability as there was no significant change of EPR signal intensity with time.
The raw data behind the stability plot is given below, in Table S1
Table S1: Spin-count data for EPR stability analysis of three replicate samples each of irradiated alanine dispersed in glycerol, measured over time. This is plotted graphically in Fig. 2B.
Week 30 kGy 50 kGy 70 kGy
0 3340 2910 3550 5590 4100 3370 6210 5420 4970
2 5070 2110 2800 5590 4740 4000 5070 4070 5460
4 2120 4030 4580 4990 6190 3740 5900 5390 4740
16 2780 2130 1890 4300 2360 3100 5100 3760 5040
6.5 Polarisation characterisation
As summarised in Tab. S2, we found that an approximately inverse relationship between build-up time constants and irradiation dose, i.e. the greater the radical concentration, the longer the slow component of the build-up time and the shorter the fast component.
This is potentially because of the competing effects of increased polarisation transfer efficiency and enhanced nuclear relaxation at higher radical concentrations. The shortened fast component likely reflects more efficient initial polarisation transfer due to increased electron-nuclear spin coupling, while the extended slow component may indicate limitations in long-range spin diffusion as paramagnetic centres become more densely packed. At higher radical concentrations, the shortened inter-radical distances could impede the cooperative network of nuclear spins necessary for efficient polarisation distribution throughout the sample, creating isolated ‘islands’ of polarisation around each radical centre. This spatial heterogeneity in the polarisation landscape would manifest as different characteristic time constants in the bi-exponential build-up curve.
At very high radical concentrations, it may be the case that the electronic environment in the crystal is similar to that of a doped semiconductor with a complex band-structure, at which point existing theories of DNP are likely not complete.
Table S2: Where $T$ is the fast component of the build-up time and $t$ is the slow component of the build-up time.
Radiation dose (kGy) Build-up Amp: P (a.u.) $T$ (s) $t$ (hours)
30 1520000 22.9 0.92
50 2280000 22.2 2.69
70 5110000 14.9 6.88
6.6 Imaging experiment details
The EPSI sequence used is described in detail elsewhere (?) and was intended to both show the spatial localisation of the hyperpolarised contrast agent and additionally be sensitive (i.e. able to resolve) any labelled impurity peaks that may be present. The acquisition matrix size was 16 $\times$ 16 (spatial) $\times$ 256 (spectral); in an $\times80$ mm2 field of view; slice thickness: 30 mm; spectral bandwidth: 5000 Hz; for a spectral resolution of 19.5 Hz/point prior to reconstruction. Data were reconstructed with zerofilling by a factor of two in all dimensions. The TR was 16.0 seconds; TE: 2.642 ms; with a 15 flip angle. Eight images were acquired.
6.7 EPR Supplementary Methods
Quantitative EPR spectra were recorded on an X-band continuous wave (CW) EPR spectrometer (EMXmicro, Bruker Biospin GMBH) using an ER 4122SHEQ resonator at room temperature; low temperature measurements, including electronic relaxation constants $T_{1e}$ and measured $T_{2e}$ (known as $T_{m}$) measurements were performed at 5 K on an X-band pulsed E680 spectrometer. Magnesium Oxide (MgO) powder, a common standard reference in EPR, (?) was used to account for any field offset in the experimental data, arising on physical grounds, which was calculated at $\sim 0.6$ mT. Furthermore, it is known that the signal amplitude is proportional to $\chi_{m}Q\sqrt{\text{Microwave power}}$ in EPR and these parameters were controlled appropriately. (?)
Samples that comprised dry powder were directly loaded into 4 mm thin wall quartz EPR tubes and their mass recorded. For repeated measurements, samples were rotated about their own axis by approximately 25 to 35  to ensure adequate powder averaging. Liquid state samples were first loaded in 2 mm diameter capillary tubes which in turn were loaded into the EPR tubes.
All solid state samples used the standard set of parameters listed below where only frequency changed between samples. The frequency has been normalised to for all data sets presented. Scans other than at 5 K were performed at room temperature. In brief, spectrometer parameters were: microwave attenuation = 36 dB; microwave power = 0.05024 mW; scans = 1; centre field = 333.8 mT; sweep width = 33 mT; sweep direction = up; sweep time = 40.04 s; time constant = 20.48 ms; conversion time = 14.46 ms; modulation amplitude = 1.2 G; modulation Frequency = 100 kHz; sweep rate = 0.816 mT/s. Echo Detected field sweeps for gamma irradiated alanine were recorded at 5 K using a standard Hahn echo pulse sequence ($\frac{\pi}{2}-\tau-\pi-\tau-\text{detect}$) with $\pi/2$ pulse length of 16 ns, $\pi$ pulse length of 32 ns and $\tau=300$ ns, a shot repetition rate of 1.024 s was used corresponding to a sweep rate of 0.2 mT/s with 1 shot per point and one average used. For ease of comparison to CW data this field swept data is presented in the derivative form in the main text. $T_{1e}$ was measured by fitting to a functional form $y=c-ae^{-\frac{x}{b}}$, where $b$ is the $T_{1}$ time, using the “fit” function in Matlab, performing a least-squares fit. The field sweep was for the sample containing OX063 was measured using free induction decay detection with a 1000 ns pulse followed by a detection period of 600 ns with a shot repetition rate of 0.2 s with 2 shots per point and a sweep width of 3 mT corresponding a sweep rate of 0.06 mT/s. Considering a g-value of 2.00319 it was found that this system had a field offset of 0.624 mT and spectra have been corrected accordingly.
For longitudinal studies observing the stability of radical species, EPR spin counts were recorded on the same day as their irradiation and then two weeks, four weeks, and 16 weeks post irradiation. Dry powder samples were prepared directly in the EPR tubes and rotated about their own axis as above; the mass of empty and filled tubes were recorded and monitored over time and data are corrected for the mass of samples present. Solutions containing glycerol were prepared as stock solutions which were used to fill haematocrit tubes which themselves were inserted into the EPR tubes. Best practices for undertaking quantitative EPR spectroscopy were followed, such as the use of an appropriate resonator, and reproducible packing metrics with impurity controlled materials. (?, ?) EPR data was analysed in Matlab. Spin counts were determined by double numerical integration of the EPR spectra. First, the raw spectra were baseline-corrected using a first-order correction, followed by integration to obtain absorption spectra. Q-value normalisation was performed to that using a reference Q-value of 8000 (obtained with the MgO sample) to account for variations in cavity loading. A polynomial fit (order 15) was then applied to specific field regions (317.5–333 mT and 337–350 mT) for secondary baseline correction. The final double integration value was calculated from the fully corrected absorption spectrum, with verification performed by comparing the numerical derivative of the absorption spectrum against the original intensity data. Full width at half maximum (FWHM) values were also calculated for each spectrum to characterise line broadening effects related to radical concentration and interactions.
For conversion between spin counts and radical concentrations, a calibration curve was established using TEMPO samples of known concentration, allowing quantitative determination of radical concentration in alanine samples.
The $T_{1e}$ for gamma irradiated alanine at 5 K recorded at the center of the observed signal was measured using an inversion recovery experiment with the sequence $\pi-t-\frac{\pi}{2}-\tau-\pi-\tau-\text{detect}$, with $\pi/2$ pulse length of 16 ns, $\pi$ pulse length of 32 ns and $\tau=$ 300 ns. The measured $T_{1}$ was greater than the allowed maximum shot repetition rate of the spectrometer and thus the command tsleep was used to generate the inter pulse delay t which gave an effective shot repetition rate of 60 s. The measured value of $T_{1e}$ for the SAR at 5 K in glycerol was 7.59 s, and $T_{m}\approx 332~ns$. For comparison, a “literature prep” (non-irradiated [1-13C]L-alanine, 18.94M NaOH, DMSO, 15 mM OX063 and 0.3 mM Dotarem gadolinium as Gd(III)) had $T_{1e}=39~ms$ (using a standard inversion recovery with a shot repetition rate of 0.25 s), Tm=378nsTm=378~ns. A “literature prep without gadolinium”, i.e. just OX063, using a shot repetition rate of 9 s and tsleep had $T_{1e}=1.41~s$ and $T_{m}=370~ns$. These fits are shown below, in Fig. S4. Raw data is provided as further SI to this manuscript.
Figure S4: Electronic $T_{1e}$ and $T_{m}$ measurement data and fits, obtained at 5 K, for A: irradiated 100 kGy alanine; B: a “literature prep” trityl radical without gadolinium; and C: a “literature prep with gadolinium” sample.
[…]
Figure S5: Spin counts recorded from 13C-alanine (black crosses, n=3) were consistent with those recorded from NA alanine (blue hexagons, n=4) at irradiation doses 30, 50, 70, and 100 kGy.
6.8 Alanine SAR EPR Fitting and EPR Simulation Details
To try to fit the EPR data obtained of irradiated powder 13C alanine, a detailed simulation was undertaken in EasySpin using appropriate previously published hyperfine constants and $g$-tensors for the radical R1 only, which is responsible for the majority of the signal recorded via EPR. This was motivated by a desire to have a single radical species for further large-scale spin-density simulations (although we note that a semi-empirical approach was additionally undertaken, using purely the experimental EPR spectrum scaled for field). The initial fitting model included a spin system with an $S=1/2$ electron and a $g$-tensor consistent with the R1 alanine radical ($g=[2.0041,2.0034,2.0024]$). The model incorporated hyperfine couplings for five protons (four $\beta$-protons and one $\alpha$-proton) with starting values derived from NA alanine, plus the 13C nucleus . These are stated in full in the attached analysis code in Data S2. We employed the pepper function provided by EasySpin (?) with a hybrid method for orientational averaging, using an adaptive grid approach with an initial grid size of [10 2] and a minimum grid size of 2. This was progressively refined throughout the optimisation process. Isotropic line broadening from unresolved interactions was modelled with a peak-to-peak linewidth (lwpp) parameter of 0.45 mT.
The experimental spectra were first-derivative spectra collected at an X-band microwave frequency of approximately 9.4 GHz (with the exact value extracted from experimental parameters) with a field range covering 330-350 mT. Spectra were baseline-corrected prior to fitting by subtracting the mean of the first 128 points.
Fitting was performed using a two-stage optimisation approach. First, we employed Matlab’s surrogateopt function with 10,000 maximum function evaluations, exploring a bounded parameter space centred around initial estimates. The objective function combined both spectral and integral matching metrics to ensure proper reproduction of both line shapes and relative intensities. Specifically, we minimised:
$\min_{\theta}\|\mathcal{I}_{\text{fit}}(\theta)-\mathcal{I}_{\text{exp}}\|_{2}+\|\mathcal{S}_{\text{fit}}(\theta)-\mathcal{S}_{\text{exp}}\|_{2}$ (S6)
where $\mathcal{I}$ represents cumulative integrals of spectra and $\mathcal{S}$ represents the normalised spectra themselves.
The final fit was further refined using Bayesian optimization (bayesopt) with additional iterations (150) to ensure convergence to the global minimum. The fitting achieved satisfactory spectral reproduction with the dominant R1 species, though slight deviations suggest potential minor contributions from R2 and R3 species or other structural heterogeneities in the 13C-labelled samples. Full experimental metadata, simulation code, and raw experimental data are provided as supplemental data S2.
We should note that the alanine radical systems formed following irradiation by ionising radiation have been studied in depth by EPR, and many reported parameters have been obtained for both the initial high-energy primary radical, and three subsequent radical species. A reasonable fit to our data can be obtained by previously reported parameters, (?) shown in Fig. S6, and these are enclosed (with EPR relaxometry data) as Data S2. These fits do not include any degree of crystallographic ordering, as rotating the tube did not affect signal intensity. At low temperature, it is likely that radical species interconvert and in particular may lose direct molecular character, with some degree of between-molecule interactions possible within the molecular sites that are not equivalent in the unit cell. The low temperature data can be modelled to a moderate level using parameters derived from those published for the primary radical, (?) summed with smaller contributions from the known SAR species R1, R2, R3. (?, ?) See code in Data S2 for specific values.
With regards to partial ordering which is invoked in the analysis of the XRD and DNP datasets but not explicitly observed in the EPR data, we should further stress the degree to which sample preparation may affect this parameter. Both DNP and XRD samples are mildly compressed into their containers, either manually when entering the sample chamber (XRD) or by 3 bar helium gas as part of an atmospheric purge cycle (DNP). In contrast, EPR tubes are narrow diameter and it was not readily possible to pack powder samples densely within them, which again might mean that EPR undertaken in these spectrometers is not representative of that within the DNP cryostat. Furthermore, reflecting the presence of different interaction strengths at different field and temperature combinations, the exact microscopic picture may well depend upon parameters such as the rate at which the sample was lowered and frozen in the magnetic field in the DNP polariser. Furthermore, the ordering we report in both the crystal and XRD instrument frames may well be aligned differently to that in our DNP cryostat, itself with $B$ aligned perpendicular to the laboratory floor (unlike EPR spectrometers, where it is parallel). Although the crystallite size distribution from XRD data is consistent with microcrystallites having a narrow size distribution of mean length $580.8\pm 37.6$ nm, and therefore the effect of gravity is expected to be minimal, such considerations cannot be ruled out as influencing the distribution of environments contained within the sample.
Figure S6: A demonstration of simulated literature values for reported alanine radicals. At room temperature, a moderately good fit to (A) natural abundance data can be obtained from a 70%/30% mixture of the radicals SAR1 and SAR2. A similar fit can be obtained for (B) the 13-C labelled alanine powder. At low temperature (C), a freely fitted mixture of R1, R2 and the primary radical can be found to approximate the shape of the experimental data. These data have been field corrected and frequency corrected to 9.6 GHz
6.9 X-ray diffraction texture analysis
Figure S7: A comparison of (1) the algebraic sum of alanine power and separately acquired glycerol XRD data (blue), and (2) the separately acquired spectrum from an alanine powder + glycerol mixture (orange). A clear change in crystallographic texture is evident from the variation in the relative intensity of the Bragg diffraction peaks. The presence of these peaks indicates that the alanine crystallites have not dissolved in glycerol.
High-quality diffraction data was obtained with a signal-to-noise ratio of approximately $10^{12}$ with visually clear differences between the two samples, irrespective of the higher (expected) polynomial background due to excess liquid glycerol in the case of the sample with glycerol (Fig. S7). Quantitative Rietveld refinement was performed using MAUD (Materials Analysis Using Diffraction, version 2.9993), a validated Java-based crystallographic software. Initially, the known orthorhombic crystal structure of L-alanine was employed as a starting model. The following parameters were systematically refined, in order:
  1. 1.
    Background (polynomial parameters)
  2. 2.
    Sample displacement and zero-offset (-0.081  $\pm$ 0.0009 )
  3. 3.
    Unit cell parameters
  4. 4.
    Overall temperature factor (B-iso = $3.00\pm$0.15 Å2)
  5. 5.
    Instrumental profile parameters (Caglioti function for peak broadening)
  6. 6.
    Crystallographic texture
For both samples, the refinement yielded satisfactory crystallographic $R$-factors, with the glycerol-dispersed sample showing $R=10.2\%$ and weighted $R_{w}=14.6\%$, compared to $R=12.8\%$ and weighted $R=18.6\%$ for the pure powder. These values indicate good agreement between the experimental data and the structural model after accounting for texture.
Texture analysis was carried out using the E-WIMV algorithm implemented in MAUD with a spherical harmonic approach. This method involves reconstructing the orientation distribution function (ODF) from the diffraction data to quantify the degree of preferred crystallite orientation. The refinement utilised (1) an exponential harmonic texture model; (2) spherical harmonics up to order 24; (3) a regular ODF sampling grid (5  resolution); (4) multiple sample orientations as, during the scans, the samples were rotated about the azimuthal axis at a frequency of 0.5 Hz to fully characterise the crystallite distribution. The reconstructed ODF maps show clear texture and are illustrated in Fig. S8. This likely originates from the asymmetric nature of the orthorhombic unit cell.
Figure S8: Orientation distribution functions for alanine and alanine/glycerol mixtures show a clear orientation preference that is more than a factor of ten times greater than for alanine alone.
The texture is represented in multiples of random distribution (mrd), where mrd = 1 indicates random orientation, mrd >1>1 indicates preferred orientation, and mrd <1<1 indicates orientation depletion; typically mrd >3>3 or $\frac{1}{3}$ are taken as evidence of a partially ordered or ordered solid. Our analysis revealed significant texture in the glycerol-dispersed sample with maximum mrd values exceeding 10 and minimum values below 0.01 in the inverse pole plot, and more than 1600 in the ODF, indicating a highly ordered crystalline arrangement within the glycerol matrix. This contrasts with the more uniformly distributed orientations in the pure powder sample.
The pole figures were generated for the principal crystallographic directions (normal direction [ND], rolling direction [RD], and transverse direction [TD]) to visualise the orientation distribution. The texture analysis revealed that alanine crystals in glycerol adopt specific preferential orientations, likely due to interactions between the crystal faces and the glycerol matrix during sample preparation. The refined texture model was incorporated into the final Rietveld refinement, which significantly improved the fit to the experimental data, confirming that the observed intensity variations were indeed due to preferred orientation rather than structural factors. These data are provided as supplemental data S1.
6.10 Molecular Dynamics Simulations
Molecular dynamics (MD) simulations were performed to investigate the interactions between alanine and glycerol molecules, with particular focus on hydrogen bonding characteristics. Rather than starting from a crystalline alanine structure, we deliberately employed a randomly positioned initial configuration for several methodologically sound reasons: (1) to establish an upper bound on the mixing thermodynamics between alanine and glycerol, (2) to avoid kinetic trapping issues associated with the prohibitively slow crystal dissolution process (which would exceed our accessible simulation timescale of tens of ns), (3) to directly probe the inherent molecular interactions without the confounding variable of crystal lattice energy, and (4) to model potential amorphous states of alanine that may be relevant in rapid cooling scenarios such as arise when a sample is placed into the DNP hyperpolariser. This approach allowed more efficient phase space sampling while providing a reference point for potential future crystalline-based simulations.
The simulation system was constructed using the Enhanced Monte Carlo (EMC) package (v4.1.5), (?) a pre-processing module for LAMMPS. (?) A mixed system containing both alanine and glycerol molecules was generated with a total of 1000 molecules at a temperature of 300 K and pressure of 1 atm. We elected to have a smaller number of molecules with periodic boundary conditions and let the system equilibrate for a longer period of time. The initial densities were set to $1.424~g/cm^{3}$ for alanine and $1.26~g/cm^{3}$ for glycerol. The PCFF (Polymer Consistent Force Field) was employed for all simulations, as implemented in the EMC and LAMMPS packages. (?, ?) PCFF is a second-generation force field developed for organic molecules, polymers, and biomolecular systems, and has been extensively previously validated for simulating hydrogen bonding interactions in similar molecular systems, including alanine crystals. (?, ?, ?) The force field parameters were automatically generated by the EMC setup procedure. The integration timestep was set to 1 fs. Long-range electrostatic interactions were handled using the particle-particle particle-mesh (PPPM) method with a precision of 0.001. A cutoff of 1 nm was applied for both van der Waals and Coulombic interactions.
The simulation protocol consisted of the following steps:
  • System initialization: The mixed alanine-glycerol system was first constructed using EMC with random molecular placements based on the specified densities in two phases
  • Energy minimization: The system was energy-minimised to eliminate bad contacts using LAMMPS.
  • Equilibration: The system was equilibrated in two phases:
    • Initial equilibration using a Langevin thermostat at 300 K with a damping parameter of 100 fs, combined with a distance-limited NVE integration (limit of 0.1 Å) for 1000 timesteps.
    • Production equilibration in the NPT ensemble at 300 K and 1 atm pressure with temperature and pressure damping parameters of 100 fs and 1000 fs, respectively.
  • Production run: Following equilibration, a production run of 30 ns (30,000,000 timesteps) was performed in the NPT ensemble. The system state was recorded every 1000 timesteps (1 ps) for analysis. This took approximately 1 day.
Post-processing was performed in VMD, (?) and density and the radial distribution function computed at the end of the equilibration process (c.f. Fig. S9). It was found that liquid glycerol did not penetrate substantially into the alanine domain on the timescale of the simulation, suggesting limited miscibility between the two substances under these conditions. The overall radial distribution function $g(r)$ exhibited features consistent with liquid-phase behaviour, including a prominent first coordination shell peak at approximately 2 Å followed by diminishing oscillations that approached uniformity at larger distances. This $g(r)$ profile confirms the absence of long-range crystalline order and indicates a liquid-like molecular arrangement with short-range spatial correlations characteristic of hydrogen-bonded systems. While our simulation approach cannot definitively determine the global thermodynamic minimum of the system (which might involve a crystalline alanine phase separate from glycerol), it effectively establishes the relative stability of the mixed state and characterises the molecular interactions at the interface between alanine and glycerol domains. The persistence of distinct domains throughout the 30 ns simulation provides valuable information about the limited mutual solubility of these compounds under the studied conditions.
Figure S9: Molecular dynamic simulations of glycerol and alanine at 300 K. a: The initial two-phase configuration remained approximately stable throughout the 30 ns simulation, with limited penetration into molecular domains, indicating poor miscibility between the compounds; b the radial distribution function $g(r)$ for the system exhibits liquid-like characteristics with a prominent first coordination shell peak at approximately 2 Å, followed by diminishing oscillations approaching uniformity at larger distances, confirming the absence of long-range crystalline order.
6.11 Spin-dynamics simulations
In an attempt to semi-empirically predict the asymmetric frequency sweep curve, a computational study on a model of the radical-damaged singular alanine crystal was undertaken. The spatial distribution of stable alanine radicals (SARs) was modelled using the known crystallographic data from the L-alanine structure provided by neutron diffraction, assuming these are unchanged by irradiation. Since electron irradiation creates radicals at random positions within the crystal lattice, we employed bootstrap sampling (10,000 iterations) of possible radical positions to determine the statistical distribution of inter-radical distances, written using ASE in Python. (?)
[…]
For systems with multiple radical centres, simulations were extended to include two electron spins with slightly different $g$-tensors (differing by $10^{-5}$ in diagonal elements) and multiple 13C spins with coordinates derived from the crystal structure. We used a spherical tensor Liouvillian formalism. The swept microwave irradiation provided by our hardware was implemented as a frequency-modulated pulse sequence with the following parameters:
pulse-on duration $=9963~\mu\text{s}$ (S7)
pulse-off duration $=48139~\mu\text{s}$ (S8)
frequency sweep bandwidth $=25~\text{MHz}$ (S9)
modulation rate $=100~\text{Hz}$ (S10)
The frequency sweep was divided into 25 discrete slices spanning from 12.5-12.5 MHz to +12.5+12.5 MHz relative to the centre frequency. For microwave sweep simulations, we examined 10 points between 187.5 GHz and 188.5 GHz. The build-up simulations were run to a maximum time of 60 seconds, corresponding to acquired experimental data.
The powder averaging was performed using the icos-2ang-12pts grid for the basic simulations and rep-2ang-100pts-sph for more detailed models. The polarisation transfer was monitored using detection operators corresponding to the 13C $L_{z}$ Liouvillian terms. For extended simulations with multiple 13C nuclei, we incorporated 56 carbon positions from a supercell of the alanine unit cell (which contains four molecules), with the radical species positioned at sites 21 and 5 in the lattice. The electron-electron dipolar coupling was calculated from their spatial separation within the crystal structure. Periodic boundary conditions were included. The simulation ran on 200 nodes, each containing 48 core Cascade Lake processors (Intel Xeon Platinum 8268 CPU @ 2.90GHz) with 392 GB of ram, and took approximately 4 days to complete.
Figure S10: A: Simulated minimum electronic-radical distances based on a 70 kGy dose equivalent electron concentration utilising a stochastic damage algorithm in ASE. B: Predicted microwave sweep curve from Spinach using atomic coordinates from (A) with periodic boundary conditions and mimicking the experimental conditions under which DNP occurred.
6.12 Analytic models
Thermal mixing is a key mechanism in Dynamic Nuclear Polarisation (DNP), widely taken to be dominant in samples used in clinical practice, and theoretically challenging. To quantitatively assess the DNP polarisation behaviour, we implemented a numerical simulation based on Wenckebach’s framework in the low temperature limit  (?).
In this model, the electron spin system is characterised by two parameters: the inverse electron Zeeman temperature $\alpha=\hbar/k_{B}T_{Z}$ and the inverse electron non-Zeeman temperature $\beta_{NZ}=\hbar/k_{B}T_{NZ}$. This assumption splits the electronic states into those that are polarised thermally (Zeeman) and those that will ultimately transfer energy to the nuclei in the irradiation phase of DNP (non-Zeeman). This results in an electron spin polarisation as a function of resonance frequency:
$P_{S}(\omega)=\tanh\left[\frac{1}{2}\left(\omega_{0}\alpha+(\omega-\omega_{0})\beta_{NZ}\right)\right]$ (S11)
The evolution of the electron Zeeman and non-Zeeman temperatures is governed by a pair of coupled differential equations:
$F_{1}(\omega_{m},\alpha,\beta_{NZ})$ $=-2W(\omega_{m})P_{s}(\omega_{m},\alpha,\beta_{NZ})-\frac{1}{T_{1}}\left[\int g(\omega)P_{s}(\omega,\alpha,\beta_{NZ})\mathrm{d}\omega-P_{L}\right]$ (S12)
$F_{2}(\omega_{m},\alpha,\beta_{NZ})$ $=-2W(\omega_{m})(\omega_{m}-\omega_{0})P_{S}(\omega_{m},\alpha,\beta_{NZ})-\frac{1}{T_{1}}\int g(\omega)(\omega_{m}-\omega_{0})P_{s}(\omega,\alpha,\beta_{NZ})\,\mathrm{d}\omega$ (S13)
where $g(\omega)$ is the normalised ESR spectral density with centre of mass $\omega_{0}$, $\omega_{m}$ is the applied microwave frequency, $W(\omega_{m})=\frac{1}{2}\pi\omega_{1S}^{2}g(\omega_{m})$ represents the microwave power, $\omega_{1S}=\gamma_{S}|B_{1}|$ is the electron Rabi frequency, $T_{1S}$ is the electron spin-lattice relaxation time, and $P_{L}=\tanh(\frac{1}{2}\omega_{0}\beta_{L})$ is the thermal equilibrium polarisation with $\beta_{L}=\hbar/k_{B}T_{L}$ being the inverse lattice temperature. High and low temperature limits are defined by the $\omega_{0}\beta\approx 1$ boundary at roughly around 1 K.
We numerically solved this scheme in Mathematica by use of the explicit analytic calculation of its Jacobian. The function $g(\omega)$ was simulated by EasySpin under appropriate experimental conditions and a numerical interpolation function applied to extend its domain to $0,\infty]$ (defining it as zero outside of the calculated region).
These equations were evaluated using adaptive Gaussian quadrature (NIntegrate) with specified domain constraints. To ensure numerical stability, the working precision was set to 20 digits, with recursion limits extended to handle the complex structure of the integrands. This process took approximately one day on a dual Xeon gold workstation with 768 GB ram, and the simulation results are shown in Fig. S11A, with experimental integrated EPR lineshapes shown in Fig. S11B. They produce a bimodal curve qualitatively similar to that in Spinach, and do not reflect experimental results.
Figure S11: A: An analytic model of thermal mixing with the theoretical EPR lineshape given produces a bimodal build up curve. B: Integrated experimental low-temperature EPR Spectra obtained for the glycerol/alanine mixture compared to at room temperature (c.f. Fig. 2) indicated a shift in $g$.
As a semi-empirical test with the observed (and not modellable) EPR lineshape obtained, a further simulation was conducted using a semi-empirically scaled EPR lineshape based on that obtained of the alanine/glycerol mixture at 5 K. This simulation was performed in Matlab (providing an independent reimplementation of the model above) and scaled the centre linearly and width (by approximately a factor of two) of the EPR line to be that appropriate at 6.7 T under the assumption that these qualitative features would describe the system, as it is not possible for us to directly measure the EPR spectrum under DNP conditions of 1.4 K and 6.7 T. Both the high temperature and low temperature approximations as defined by Wenckebach have been solved separately and compared to experimental data. Even at 30 kGy irradiation, where a small amount of bimodal behaviour is observed, this model (which accurately reproduces the behaviour of e.g. trityl radicals, TEMPO, Totapol, and other samples) does not fit the data obtained empirically, shown in Fig. S12. We are confident that the bandwidth of the (commercially supplied) microwave excitation source exceeds that shown, and is not the limiting factor in these experiments.
Figure S12: A semi-empirical model of the absolute value of the predicted enhancement curve of DNP as a function of frequency. The EPR spectrum used in this prediction was that obtained at 5 K and lower field, scaled approximately for the system at hand. Whilst some degree of asymmetry is predicted by the Wenckebach model, it is not observed. An empirical shift of 90 MHz aligns the curves, but does not resolve the predicted asymmetry.
6.13 QuantumESPRESSO
First-principles electronic structure calculations were performed using the plane-wave pseudopotential approach as implemented in the Quantum ESPRESSO (QE) package (version 7.2)  (?, ?). The Perdew-Burke-Ernzerhof (PBE) generalised gradient approximation  (?) was used for the exchange-correlation functional. Projector augmented-wave (PAW) pseudopotentials were employed for carbon and oxygen atoms, while ultrasoft pseudopotentials were used for hydrogen and nitrogen atoms, all obtained from the Standard Solid-State Pseudopotential (SSSP) library (efficiency version)  (?).
The crystal structure of L-alanine was based on the available neutron diffraction data, with a unit cell containing 52 atoms (four alanine molecules). The orthorhombic unit cell parameters were $a=5.7880$ Å, $b=6.0360$ Å, and $c=12.3420$ Å, with space group P212121 (No. 19).
Calculations were performed with the following parameters: kinetic energy cutoff for wavefunctions of 60 Ry; kinetic energy cutoff for charge density of 480 Ry; 100 Kohn-Sham bands computed to ensure proper description of the conduction bands; electronic convergence threshold of $1.0\times 10^{-6}$ Ry; mixing parameter (beta) of 0.6.
Brillouin Zone Sampling
For the band structure calculations, a path along high-symmetry points in the first Brillouin zone was sampled. A total of 392 k-points were used along the path, which included the following segments: $\Gamma\rightarrow$ X; X $\rightarrow$ S; S $\rightarrow$ Y; Y $\rightarrow\Gamma$; $\Gamma\rightarrow$ Z; Z $\rightarrow$ U; U $\rightarrow$ R; R $\rightarrow$ T.
The k-point mesh was generated in crystal coordinates, with a dense sampling to ensure smooth band dispersion.
Computational Resources
The calculations were performed on a parallel computing architecture using 40 MPI processes distributed across a single node and using nearly 1 TB of ram. The electronic structure calculation for the system consumed approximately 24 hours of CPU time and 31 hours of wall time, highlighting the computational intensity of modelling this organic crystal system.
Analysis of Electronic Structure
The band structure calculations revealed a fundamental electronic band gap of approximately 5.02 eV (calculated as the difference between the highest occupied state at 0.58 eV and the lowest unoccupied state at 5.60 eV). The energy eigenvalues were extracted from the full band structure calculation and used to generate the band dispersion plots shown in Fig. S13. This compares favourably to one reported experimental values of 5.4 eV, (?) especially as DFT studies using PBE functionals tend to underestimate band gaps by 0.5-1.0 eV.
The total energy of the system was 668.849-668.849 Ry, with contributions from: one-electron term (362.629-362.629 Ry); Hartree term (237.420 Ry); exchange-correlation term (157.484-157.484 Ry); Ewald term (236.071-236.071 Ry); one-center PAW contribution (150.084-150.084 Ry).
Figure S13: The estimated bandstructure of alanine crystals predicts that it is a relatively large bandgap insulator.
The relatively large bandgap supports the stability of radical environments, and rules out one (admittedly outlandish) potential mechanism for DNP: that of the well-resolved solid effect.
6.14 Products of dissolution
There were no 13C-labelled impurity peaks present in the dissolution products, and the produced hyperpolarised spectra were largely indistinguishable from that using trityl radicals. This is shown in detail below, in Fig. S14.
Figure S14: Hyperpolarised alanine produced either with trityl radical or with e-beam irradiation, both when initially dissolved (1 average, WALTS-16 proton decoupling) and corresponding thermal-equilibrium acquisition (256 averages, fully relaxed, WALTZ-16 proton decoupling). No impurity peaks were observed in either spectra. Data obtained at 1.4 T on a Magritek Spinsolve Ultra. The enhancement factor is approximately 142,000.
7 Data captions
Caption for Data S1.
Raw and analysed crystallographic structures for Alanine are provided as Data S1, including structural CIF files and textural fits and refinements used in this work.
Caption for Data S2.
Raw EPR data of irradiated and natural abundance alanine used for modelling in this work, together with MATLAB files that reproduce either the modelled spectrum or $T_{1e}$ or $T_{m}$ fits.